### v1

# algorithmic

Abstract
In this paper, we study the problem of efficient online reinforcement learning in the infinite horizon setting when there is an offline dataset to start with. We assume that the offline dataset is generated by an expert but with unknown level of competence, i.e., it is not perfect and not necessarily using the optimal policy. We show that if the learning agent models the behavioral policy (parameterized by a competence parameter) used by the expert, it can do substantially better in terms of minimizing cumulative regret, than if it doesn’t do that. We establish an upper bound on regret of the exact informed PSRL algorithm that scales as $\tilde{O}(\sqrt{T})$. This requires a novel prior-dependent regret analysis of Bayesian online learning algorithms for the infinite horizon setting. We then propose an approximate Informed RLSVI algorithm that we can interpret as performing imitation learning with the offline dataset, and then performing online learning.
1 Introduction
The original vision of Reinforcement Learning (RL) was that of a learning agent taking actions, getting state and reward feedback and learning from it. The learning agent may have some prior information but learning mostly happens from these online interactions. This requires carefully-crafted strategies that balance exploration and exploitation, including the need for deep exploration in many sequential decision-making problems. This was regarded as a difficult problem for a long time. Fortunately, over the last decade or so, substantial progress has been made on the online reinforcement learning (ORL) problem. Unfortunately, ORL Algorithms tend to be rather “data-hungry”, i.e., they require a lot more interaction data than other types of learning methods.
In some ORL problem settings, offline datasets are available. There is thus a possibility of improving learning efficiency by using offline datasets for pre-training, and then the learning agent does further fine-tuning upon deployment for online learning. For example, such offline could be demonstrations from an expert, or even a sub-expert. This allows for addressing the distribution shift problem with the offline reinforcement learning methods, i.e., policy learnt offline do not perform well because the distribution of the system upon deployment is different from the one encountered offline. One can also expect that when such offline datasets are available for pre-training, it would improve online learning efficiency (in terms of cumulative regret). The question is by how much, and how does it depend on the dataset quality, and indeed how do we even quantify dataset quality. Furthermore, even with the same dataset, a question arises on how to extract the most from such a given dataset.
An important reason via a combination of offline and online learning is important is that reinforcement learning from offline datasets can often suffer from the distribution shift problem, i.e. the distribution of the system upon deployment is different from the one encountered offline, and hence the policy learnt offline do not perform well upon deployment. Learning from offline dataset first followed by online learning allows for adaptation of learnt policies to different distributions of the underlying environment.
In this paper, we develop a systematic method to incorporate expert demonstrations into online learning algorithms to facilitate faster learning for infinite horizon MDPs. We first propose the idealized infinite-horizon informed Posterior Sampling based Reinforcement Learning (inf-iPSRL) algorithm. Under some mild assumptions on the underlying MDP, we show that the algorithm achieves $O(1)+\tilde{O}(\sqrt{\varepsilon T})$ where $\varepsilon$ is the estimation error probability of the optimal policy given the offline data, and $T$ is the learning horizon. The regret bound also has polynomial dependence on problem parameters of the underlying MDP. We also show that, if the offline data is generated by an expert with suitably high competence level, under certain assumptions, the regret of the inf-iPSRL algorithm goes to $O(1)$ as the size of the demonstration dataset goes to infinity. We would also like to point out the inf-iPSRL algorithm has a simpler design compared to the TSDE algorithm in Ouyang et al. (2017), since the TSDE algorithm uses a visitation-count-based episode schedule while the inf-iPSRL algorithm uses fixed episode schedules. However, the inf-iPSRL algorithm is still impractical due to the complicated nature of exact posterior updates. Therefore we introduce the informed randomized least-squares value iteration (inf-iRLSVI) algorithm which replaces exact posterior sampling with an approximation procedure. Just like the RSVI algorithm of Osband et al. (2019) (from which we borrowed the name), the inf-iRLSVI algorithm generates approximate posterior samples via optimizing a randomly perturbed loss function.
Related Work. Recently, offline reinforcement learning, where offline datasets are incorporated (Levine et al., 2020), has attracted some interest due to the widespread practice of pre-training with offline datasets in large language models (Brown et al., 2020; Thoppilan et al., 2022; Hoffmann et al., 2022). One fundamental challenge in this line of research is the distribution shift problem, where the policy learnt from offline data performs poorly in the real world. To address this problem, Jin et al. (2021); Rashidinejad et al. (2021); Xie et al. (2021a) adopted the pessimistic approach, which can be overly conservative in practice. In Uehara and Sun (2021); Rashidinejad et al. (2021); Xie et al. (2021a); Agarwal and Zhang (2022), the authors identified sufficient conditions on the dataset that guarantees a certain level of performance for offline RL algorithms. However, determining whether the dataset meets these conditions can be a complicated problem itself, hence making these results impractical (Kumar et al., 2020; Nair et al., 2020; Argenson and Dulac-Arnold, 2020; Levine et al., 2020; Kostrikov et al., 2021; Wagenmaker and Pacchiano, 2022).
On the other hand, there have been many works on online reinforcement learning (See Russo et al. (2018) for a survey). Among these works, there are two prominent approaches: the optimism under the face of uncertainty (OFU) approach (Auer et al., 2008) and posterior sampling (PS) approach (Osband et al., 2013; Russo and Van Roy, 2016; Ouyang et al., 2017). Most work in online RL focuses on the online dataset and does not consider leveraging the offline dataset. Closely related to the idea of using the offline dataset is the concept of imitation learning (Schaal, 1996; Hester et al., 2018; Beliaev et al., 2022), where one aims to learn the behavioral policy of the expert from the offline dataset. No online finetuning is present in these works. In Ernst et al. (2005); Vecerik et al. (2017); Rashidinejad et al. (2021); Hansen et al. (2022); Kumar et al. (2022); Lee et al. (2022), the authors combine imitation learning with more traditional offline RL methods. In (Schrittwieser et al., 2021; Uehara and Sun, 2021; Xie et al., 2021b; Agarwal and Zhang, 2022; Song et al., 2022; Fang et al., 2022; Wan et al., 2022; Ball et al., 2023), the authors combined offline RL with limited online policy fine-tuning to minimize the simple regret.
2 Preliminaries
We first formally introduce our model, assumptions and problem statement.
We assume that a learning agent interacts with a time-homogeneous controlled Markov chain over time, where an instantaneous reward is generated at each time. The process is represented by a tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},P,r,\nu)$, where $\mathcal{S}$ is the state space; $\mathcal{A}$ is the action space; $P:\mathcal{S}\times\mathcal{A}\mapsto\Delta(\mathcal{S})$ is the transition kernel (where $\Delta(\mathcal{S})$ represents the set of distributions on $\mathcal{S}$); $r:\mathcal{S}\times\mathcal{A}\mapsto[0,1]$ is the instantaneous reward function; $\nu\in\Delta(\mathcal{S})$ is the distribution of the initial state. At time $t=0$, the initial state $s_{0}$ is drawn from $\nu$. At each time $t\in\mathbb{Z}_{+}$, the agent chooses an action $a_{t}\in\mathcal{A}$, and the next state $s_{t+1}$ is chosen according to the distribution $P(\cdot|s_{t},a_{t})$. The objective of the agent is to choose a policy $\pi$ to maximize the expected average reward over the infinite horizon, i.e. maximizing
$\bar{r}^{\pi}:=\lim_{T\rightarrow\infty}\dfrac{1}{T}\mathbb{E}^{\pi}\left[\sum_{t=0}^{T-1}r(s_{t},a_{t})\right]$ (1)
From standard literature on MDPs (Bertsekas, 1995), we know that if the controlled Markov chain is communicating (i.e., for any pair of states $s,s^{\prime}\in\mathcal{S}$, there exists a stationary Markov policy $\pi$ and an integer $k\in\mathbb{N}$ such that $\mathbb{P}_{\theta}^{\pi}(s_{k}=s^{\prime}|s_{0}=s)>0$), the optimal expected average reward and an optimal policy can be obtained though total discounted rewards: Let $\gamma\in(0,1)$ be the discount factor. Define $V_{\gamma}^{*}(s):=\max_{\pi}\mathbb{E}^{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})|s_{0}=s\right]$ to be the optimal total discounted reward when the initial state is $s\in\mathcal{S}$. Let $\hat{s}\in\mathcal{S}$ be an arbitrary reference state, then we have
$\bar{r}^{*}$ $:=\lim_{\gamma\rightarrow 1}(1-\gamma)V_{\gamma}^{*}(\hat{s})$ (2)
to be the optimal expected average reward for this MDP with any starting state. Furthermore, an optimal policy can be recovered as follows. Define the relative state value function
$v(s)$ $:=\lim_{\gamma\rightarrow 1}[V_{\gamma}^{*}(s)-V_{\gamma}^{*}(\hat{s})]\qquad\forall s\in\mathcal{S}$ (3)
which represents the long-term difference of the optimal expected total reward if the MDP starts at $s$ comparing to starting at $\hat{s}$ (Bertsekas, 1995). Standard results show that $\bar{r}^{*}$ and $v^{*}$ satisfy the Bellman equation (Bertsekas, 1995):
$\bar{r}^{*}+v(s)=\max_{a\in\mathcal{A}}\left[r(s,a)+\sum_{s^{\prime}\in\mathcal{S}}P(s^{\prime}|s,a)v(s^{\prime})\right]$ (4)
for every $s\in\mathcal{S}$. Define the relative state-action value function
$q(s,a):=r(s,a)+\sum_{s^{\prime}\in\mathcal{S}}P(s^{\prime}|s,a)v(s^{\prime})$ (5)
then any deterministic stationary Markov policy $\pi$ such that $\pi(s)\in\operatorname*{arg\,max}_{a\in\mathcal{A}}q(s,a)$ for every $s\in\mathcal{S}$ maximizes the expected average reward.
We assume that the learning agent does not possess full knowledge about the environment $\mathcal{M}$. The agent may however, have complete or partial prior knowledge about certain parameters of $\mathcal{M}$. To provide a general model for such prior knowledge, we use $\mathcal{M}(\theta)$ to denote an environment parameterized by some parameter $\theta$ that lies in the set $\varTheta$. The true environment is given by a random variable $\theta^{*}$ with distribution $\mu_{0}\in\Delta(\varTheta)$. We make the following assumptions on the environment.
Assumption 2.1.
For all $\theta\in\varTheta$, $\mathcal{M}(\theta)$ is a communicating MDP.
In a general infinite horizon MDP, it is possible that one decision made at a certain time can have an irreversible effect on the total reward: It is possible that one enters a communicating class with suboptimal average reward. Assumption 2.1 ensures that one could explore the MDP using any policy since there’s always a pathway to return to the “optimal part” of the state space. See Ouyang et al. (2017) for more discussions.
For a communicating MDP $\mathcal{M}(\theta)$, we use $\bar{r}^{*}(\theta),v(\cdot;\theta)$, and $q(\cdot,\cdot;\theta)$ to denote their corresponding $\bar{r}^{*},v^{*},$ and $q^{*}$ functions, as defined in (2), (3) and (5) respectively. We also make Assumption 2.2, which is the same as that in Ouyang et al. (2017).
Assumption 2.2.
There exists $\bar{v}\in\mathbb{R}_{+}$ such that $\max_{s\in\mathcal{S}}v(s;\theta)-\min_{s\in\mathcal{S}}v(s;\theta)\leq\bar{v}$ for all $\theta\in\varTheta$.
Under the above assumptions we consider the canonical settings of tabular MDPs as well as linear value function generalization for non-tabular MDPs, i.e., the Q-value function is representable in a subspace spanned by a given set of features.
The learning agent has access to an offline dataset $\mathcal{D}_{0}=\{(\bar{s}_{0},\bar{a}_{0},\cdots,\bar{s}_{N-1},\bar{a}_{N-1},\bar{s}_{N})\}$ generated by another agent, called the expert, who interacts with the same environment before the learning agent started. The trajectory is generated from applying a randomized stationary Markov policy on the true MDP $\mathcal{M}(\theta^{*})$. The initial state $\bar{s}_{0}$ is assumed to be a fixed, non-random state. We will describe the expert’s policy and the learning agent’s knowledge about it later in this section. During the learning process, the learning agent has access to the trajectory generated so far, which we call the online dataset and denote as $\mathcal{H}_{t}=\{(s_{0},a_{0},\cdots,s_{t-1},a_{t-1},s_{t})\}$. At time $t$, the learning agent’s information is given by $\mathcal{D}_{t}:=\mathcal{D}_{0}\cup\mathcal{H}_{t}$.
The learning agent utilizes a learning algorithm $\phi$ to interact with the environment, which at each time $t$ chooses a possibly randomized action based on full or partial information in the dataset $\mathcal{D}_{t}$. Under the environment $\mathcal{M}(\theta)$, the learning regret is defined as
$\mathfrak{R}_{T}(\phi;\theta):=T\bar{r}^{*}(\theta)-\mathbb{E}_{\theta}^{\phi}\left[\sum_{t=0}^{T-1}r_{t}(s_{t},a_{t})\right],$ (6)
where $\mathbb{E}_{\theta}^{\phi}$ denotes the expectation under the possible trajectories generated by applying algorithm $\phi$ on the environment $\mathcal{M}(\theta)$. The Bayesian regret is defined as
$\mathfrak{BR}_{T}(\phi):=\mathbb{E}_{\theta^{*}\sim\mu_{0}}[\mathfrak{R}_{T}(\phi;\theta^{*})]$ (7)
We model the expert’s policy in the following way. First, consider that the expert, who generated the offline dataset, has perfect knowledge about the underlying environment $\theta^{*}$ but uses an approximately optimal policy in the demonstration. Then, the expert’s policy $\pi^{\beta}$ is modelled as:
$\pi^{\beta}(a|s):=\dfrac{\exp(\beta(s)q(s,a;\theta^{*}))}{\sum_{a^{\prime}\in\mathcal{A}}\exp(\beta(s)q(s,a^{\prime};\theta^{*}))}$ (8)
where $\beta(s)\geq 0$ is called the state-dependent deliberateness parameter: $\beta(s)=0$ means that the expert takes actions uniformly randomly at state $s$; A large $\beta(s)>0$ means that the expert uses a close-to-optimal policy at state $s$.
Then, we consider a general model wherein the expert only knows a noisy version of the optimal $q$-function. The expert uses the policy (8), except that $q^{*}$ is replaced by $\tilde{q}$, the expert’s perceived state-action value function. The vector $\tilde{q}$ follows Gaussian distribution with mean $q^{*}$ and covariance $\mathbb{I}/\lambda^{2}$, where $\lambda>0$ is the knowledgeability parameter. The two parameters $(\beta,\lambda)$ together will be referred to as the competence of the expert. In this case, we denote the expert’s (stationary) policy as $\pi^{\beta,\lambda}$. Such a model was first introduced in Hao et al. (2023).
The learning agent knows the form of the expert’s policy and the knowledgeability parameter $\lambda$ but may not know the deliberateness parameter $\beta$. When $\beta$ is unknown to the learning agent, we will assume an independent exponential prior for the sake of analytical simplicity, $f_{2}(\beta(s))=\lambda_{2}\exp(-\lambda_{2}\beta(s))$ over $\beta(s)$ where $\lambda_{2}>0$ is the same for all $s$.
3 The Infinite-horizon Informed PSRL Algorithm for Average MDPs
We now introduce the Informed Posterior Sampling-based Reinforcement Learning (inf-iPSRL) algorithm (Algorithm 1) for infinite-horizon average MDPs that combines the offline data with the powerful posterior sampling method (Osband et al., 2013; Ouyang et al., 2017). We assume that the expert knows the exact $q(\cdot,\cdot;\theta^{*})$ function and follows the policy $\pi^{\beta}$ given by (8) with $\beta(s)=\beta>0$ for all $s\in\mathcal{S}$. We assume that the learning agent knows $\beta$ but not $\theta^{*}$. Therefore, the posterior distribution of $\theta^{*}$ given the offline data $\mathcal{D}_{0}$ is given by
$\mu_{1}(\theta^{*}\in\cdot)\propto\int_{\theta\in\cdot}\mathbb{P}_{\theta}^{\pi^{\beta}}(\mathcal{D}_{0})\mathrm{d}\mu_{0}(\theta),$ (9)
where
$\mathbb{P}_{\theta}^{\pi^{\beta}}(\mathcal{D}_{0}):=\prod_{t=0}^{N-1}\theta(\bar{s}_{t+1}|\bar{s}_{t},\bar{a}_{t})\dfrac{\exp(\beta q(\bar{s}_{t},\bar{a}_{t};\theta))}{\sum_{a^{\prime}\in\mathcal{A}}\exp(\beta q(\bar{s}_{t},a^{\prime};\theta))}.$ (10)
In the inf-iPSRL algorithm we assume the availability of an planning oracle, called MDPSolve, to output an optimal policy for the MDP $\mathcal{M}(\theta)$ given input $\theta$: We assume that MDPSolve is a deterministic mapping from the space of parameters $\varTheta$ to the space of deterministic stationary Markov strategies $\Pi$ with the following property: $\pi=\texttt{MDPSolve}(\theta)$ satisfies $\pi(s)\in\operatorname*{arg\,max}_{a\in\mathcal{A}}q(s,a;\theta)$ for all $s\in\mathcal{S}$. Note that various dynamic programming algorithms such as value iteration, policy iteration, etc. can be used in MDPSolve.
The inf-iPSRL algorithm divides the learning process into episodes, with possibly different lengths. The episode lengths $\{T_{k}\}_{k=1}^{K}$ are fixed and given as an input to the algorithm. Let $t_{k}:=\sum_{j=1}^{k-1}T_{k}$ denote the starting time of episode $k$. At the beginning of episode $k$, the learning agent randomly samples a new environment $\tilde{\theta}^{k}$, uses the planning oracle MDPSolve to find an average-reward maximizing policy $\tilde{\pi}^{k}$ for $\tilde{\theta}^{k}$, and apply the policy $\tilde{\pi}^{k}$ for the duration of episode $k$. Finally, after applying $\tilde{\pi}^{k}$ for $T_{k}$ steps, the learning agent updates the posterior distribution of $\theta^{*}$ via
$\mu_{k+1}(\theta^{*}\in\cdot)\propto\int_{\theta\in\cdot}\prod_{t=t_{k}}^{t_{k+1}-1}\theta(s_{t+1}|s_{t},a_{t})\mathrm{d}\mu_{k}(\theta).$ (11)
Algorithm 1 inf-iPSRL Algorithm
  Input: Prior $\mu^{0}\in\Delta(\varTheta)$; Offline data $\mathcal{D}_{0}$; Episode schedule $\{T_{k}\}_{k=1}^{K}$
  $t_{1}\leftarrow 0$
  Compute $\mu_{1}$ with (9)
  for $k=1$ to $K$ do
     Sample $\tilde{\theta}^{k}\sim\mu^{k}$
     $\tilde{\pi}^{k}\leftarrow\texttt{MDPSolve}(\tilde{\theta}^{k})$
     $t_{k+1}\leftarrow t_{k}+T_{k}$
     for $t=t_{k}$ to $t_{k+1}-1$ do
        Take action $a_{t}=\tilde{\pi}^{k}(s_{t})$
     end for
     Compute $\mu_{k+1}$ with (10)
  end for
3.1 Prior Dependent Bound
In this section, we develop a prior-dependent regret bound for the inf-iPSRL algorithm on the tabular RL setting. The regret bound depends on the prior distribution $\mu_{1}$ through the initial mismatch probability $\varepsilon=\mathbb{P}(\tilde{\pi}^{1}\neq\pi^{*})$, where $\pi^{*}=\texttt{MDPSolve}(\theta^{*})$ is the optimal policy of the underlying true MDP consistent with planning oracle we used in the inf-iPSRL algorithm. Intuitively, if the offline data is more informative, then the initial posterior distribution on the optimal policy $\pi^{*}$ is more concentrated, and then it’s more likely that the initial policy obtained from posterior sampling is actually the optimal policy, meaning that $\varepsilon$ is small. We will formalize this at the end of this section.
Theorem 3.1.
Let $K_{T}:=\max\{k:\sum_{l=1}^{k-1}T_{k}<T\}$. The Bayesian regret for inf-iPSRL algorithm satisfies
$\mathfrak{BR}_{T}(\phi^{\mathrm{iPSRL}})\leq 3\bar{v}+2\bar{v}(R_{1}+R_{2}+R_{3}),$ (12)
where
$R_{1}$ $:=\varepsilon K_{T},$ (13)
$R_{2}$ $:=\sqrt{\varepsilon S^{2}AT\log(2SAK_{T}T)\left[1+\log\left(\frac{T}{SA}+1\right)\right]},$ (14)
$R_{3}$ $:=\min\left\{\varepsilon T,SA\left(\max_{k\in[K_{T}]}T_{k}\right)\log_{2}\left(\frac{T}{SA}+1\right)\right\}.$ (15)
Note that the right-hand side of (12) depends on the episode schedule only through (i) total number of episodes with time $T$, and (ii) the maximum length of an episode within time $T$. Therefore, the best regret bound, in general, can be obtained by setting both quantities to be roughly $\Theta(\sqrt{T})$, which can be achieved when $T_{k}=\Theta(k)$. We formalize this in the following two corollaries.
Corollary 3.2.
Consider the episode schedule $T_{k}=k$. Then,
$\mathfrak{BR}_{T}(\phi^{\mathrm{iPSRL}})\leq\tilde{O}(\bar{v}SA\sqrt{T}),$ (16)
where $\tilde{O}$ hides the logarithmic terms in $S,A,$ and $T$. Furthermore, for fixed $\bar{v},S,A,T$ we also have an $O(1)+O(\sqrt{\varepsilon})$ regret bound:
$\mathfrak{BR}_{T}(\phi^{\mathrm{iPSRL}})\leq 3\bar{v}+2\varepsilon\bar{v}T+\sqrt{\varepsilon}\bar{v}\cdot\tilde{O}(\sqrt{S^{2}AT}).$ (17)
Remark 3.3.
The only difference between the episode schedule of the TSDE algorithm in Ouyang et al. (2017) and inf-iPSRL algorithm with $T_{k}=k$ is a stopping criteria based on visitation count of all state-action pairs. Our result here shows that the Bayesian regret is still $\tilde{O}(\sqrt{T})$ if one removes the visitation count based stopping criteria in the TSDE algorithm.
When $\varepsilon$ is small, one can obtain a better regret bound for large $T$ by designing episode lengths accordingly.
Corollary 3.4.
Let $\hat{\varepsilon}\geq\varepsilon$. Consider the episode schedule $T_{k}=\lceil\hat{\varepsilon}k\rceil$. Then,
$\mathfrak{BR}_{T}(\phi^{\mathrm{iPSRL}})\leq 3\bar{v}+\tilde{O}(\bar{v}SA\sqrt{\hat{\varepsilon}T}).$ (18)
To better interpret the dependence of the regret on the prior distribution, we note that the initial mismatch probability can be bounded through the estimation error of the optimal policy given the offline data.
Lemma 3.5.
Let $\hat{\pi}^{*}$ be any estimator of $\pi^{*}$ constructed from $\mathcal{D}_{0}$, then $\mathbb{P}(\tilde{\pi}^{1}\neq\pi^{*})\leq 2\mathbb{P}(\hat{\pi}^{*}\neq\pi^{*})$.
Proof.
If $\tilde{\pi}^{1}\neq\pi^{*}$, then either $\hat{\pi}^{*}\neq\tilde{\pi}^{1}$ or $\hat{\pi}^{*}\neq\pi^{*}$ must be true. Conditioning on $\mathcal{D}_{0}$, $\tilde{\pi}^{1}$ is identically distributed as $\pi^{*}$ while $\hat{\pi}^{*}$ is deterministic, therefore
$\mathbb{P}(\tilde{\pi}^{1}\neq\pi^{*})\leq\mathbb{P}(\hat{\pi}^{*}\neq\tilde{\pi}^{1})+\mathbb{P}(\hat{\pi}^{*}\neq\pi^{*})=2\mathbb{P}(\hat{\pi}^{*}\neq\pi^{*})$
3.2 Bounding the Estimation Error of Optimal Policy
We now provide a bound on the estimation error of the optimal policy given the offline data. In addition to Assumption 2.1 and 2.2, we make the following additional assumptions.
Assumption 3.6.
There exists a number $\Delta>0$ such that for all $\theta\in\varTheta$ and all $s\in\mathcal{S}$, there exists $a^{*}\in\mathcal{A}$ such that $q(s,a^{*};\theta)-q(s,a;\theta)\geq\Delta$ for all $a\neq a^{*}$.
Assumption 3.7.
There exists a number $\delta>0$ such that for all $\theta\in\varTheta$ and all $s,s^{\prime}\in\mathcal{S}$, there exists a stationary Markov policy such that
$\mathbb{P}_{\theta,\pi}(\exists 0<t\leq S,s_{t}=s^{\prime}|s_{0}=s)\geq\delta.$ (19)
Here, $S$ is the number of states. Assumption 3.7 can be interpreted as requiring all MDPs to be “uniformly communicating”: If we allow $\delta$ to be dependent on $\theta$ in Assumption 3.7, then Assumption 3.7 is the same as Assumption 2.1, i.e. all $\mathcal{M}(\theta)$ are communicating.
Lemma 3.8.
For a sufficiently large $\beta(s)=\beta$ and any $N\in\mathbb{N}$, there exists an estimator $\hat{\pi}^{*}$ such that
$\mathbb{P}(\hat{\pi}^{*}\neq\pi^{*})\leq\varepsilon_{N}:=S(1+1.06N)\exp(-cN)$ (20)
where $c>0$ is a constant dependent on $(S,A,\bar{v},\Delta,\delta,\beta)$.
4 An Approximate Bayesian Algorithm
4.1 The Informed RLSVI Algorithm for average-reward MDPs
The inf-iPSRL algorithm introduced in the previous section assumes that the posterior updates can be done exactly. In practice, the posterior update in (9) is challenging due to the loss of conjugacy while using the Bayes rule. Thus, we must find a computationally efficient way to do approximate posterior updates (and obtain samples from it) to enable practical implementation. Hence, we propose a novel approach based on Bayesian bootstrapping to obtain approximate posterior samples. The key idea is to perturb the loss function for the maximum a posterior (MAP) estimate and use the point estimate as a surrogate for the exact posterior sample.
Note that in the ensuing, we regard $\beta$ as also unknown to the learning agent (and $\lambda=\infty$ for simplicity). Thus, the learning agent must form a belief over both $\theta$ and $\beta$ via a joint posterior distribution conditioned on the offline dataset $\mathcal{D}_{0}$ and the online data at time $t$, $\mathcal{H}_{t}$. We denote the prior pdf over $\theta$ by $f(\cdot)$ and prior pdf over $\beta$ by $f_{2}(\cdot)$.
Consider the average-reward Bellman optimality eq.:
$\bar{r}^{*}+q^{*}(s,a;\theta)=r(s,a)+\mathbb{E}[\max_{b}q^{*}(s^{\prime},b;\theta)].$
For the sake of compact notation, we denote $q^{*}(s,a;\theta)$ as $q^{\theta}_{t}(s,a)$ in this section. Now, consider the offline dataset,
$\mathcal{D}_{0}=\{(s_{k},a_{k},\check{s}_{k},r_{k})_{k=0}^{N-1}\}$
and let $\theta$ denote the parameterization of the relative q-value function. We introduce the temporal difference error ${\mathcal{E}}_{k}$ (parameterized by a given $q^{\theta}$),
$\mathcal{E}_{k}(q^{\theta}):=\left(r_{k}+\max_{b}q^{\theta}(\check{s}_{k},b)-q^{\theta}(s_{k},a_{k})-\bar{r}_{k}\right),$
and define recursively,
$\bar{r}_{k+1}:=(1-\eta_{k})\bar{r}_{k}+\eta_{k}r_{k},~{}~{}\bar{r}_{0}:=r_{0}.$
We use this to construct a parameterized offline dataset,
$\mathcal{D}_{0}(q^{\theta})=\{(s_{k},a_{k},\check{s}_{k},\bar{r}_{k},\mathcal{E}_{k}(q^{\theta}))_{k=0}^{N-1}\}.$
A parameterized online dataset at time $t$, $\mathcal{H}_{t}(q^{\theta})$ can be similarly defined. To ease notation, we will regard the time $t$ during the online phase as the $(N+t)$ time. Thus,
$\mathcal{H}_{t}(q^{\theta})=\{(s_{k},a_{k},\check{s}_{k},\bar{r}_{k},\mathcal{E}_{k}(q^{\theta}))_{k=N}^{N+t}\}\},$
the online dataset available at time $t$.
Note that $q^{\theta}$ is to be regarded as a parameter. Now, at time $t$, we would like to obtain a MAP estimate for $(\theta,\beta)$ by solving the following:
$\begin{split}\textbf{MAP:}~{}~{}~{}\arg\max_{\theta,\beta}~{}&\log P(\mathcal{H}_{t}(q^{\theta})|\mathcal{D}_{0}(q^{\theta}),\theta,\beta)+\log P(\mathcal{D}_{0}(q^{\theta})|\theta,\beta)+\log f(\theta)+\log f_{2}(\beta).\end{split}$ (21)
Denote a perturbed version of the $q^{\theta}$-parameterized offline dataset by
$\tilde{\mathcal{D}}_{0}(q^{\theta})=\{(s_{k},\tilde{a}_{k},\check{s}_{k},\bar{r}_{k},\tilde{\mathcal{E}}_{k})_{k=0}^{N-1}\}$
where we add the following random perturbations: (i) actions have perturbation $w_{l}^{h}\sim\exp(1)$, (ii) rewards have perturbations $z_{k}\sim\mathcal{N}(0,\sigma^{2})$, and (iii) the prior $\tilde{\theta}\sim\mathcal{N}(0,\Sigma_{0})$.
Note that the first and second terms involving $\mathcal{H}_{t}$ and $\mathcal{D}_{0}$ in (21) are independent of $\beta$ when conditioned on the actions. Thus, we have a sum of log-likelihood of TD error, transition and action as follows:
$\begin{split}&\quad~{}\log P(\tilde{\mathcal{D}}_{0}(q^{\theta})|q^{\theta})\\ &=\sum_{k=0}^{N-1}\left(\log P(\tilde{\mathcal{E}}_{k}|\check{s}_{k},a_{k},s_{k},\bar{r}_{k},q^{\theta})+\log P(\check{s}_{k}|a_{k},s_{k},q^{\theta})+\log P(a_{k}|s_{k},q^{\theta})\right)\\ &\leq\sum_{k=0}^{N-1}\left(\log P(\tilde{\mathcal{E}}_{k}|\check{s}_{k},a_{k},s_{k},\bar{r}_{k},q^{\theta})+\log\pi^{\beta}(a_{k}|s_{k},q^{\theta})\right).\end{split}$
By ignoring the log-likelihood of the transition term (akin to optimizing an upper bound on the negative loss function), we are actually being optimistic.
For the terms in the upper bound above, under the random perturbations assumed above, we have
$\begin{split}&\log P(\tilde{\mathcal{E}}_{k}|\check{s}_{k},a_{k},s_{k},\bar{r}_{k},q^{\theta})=-\frac{1}{2}(r_{k}+z_{k}+\max_{b}q^{\theta}(\check{s}_{k},b)-q^{\theta}(s_{k},a_{k})-\bar{r}_{k})^{2}+\text{constant}\end{split}$
and
$\begin{split}\log\pi^{\beta}(a_{k}|s_{k},q^{\theta}_{t})&=w_{k}\left(\beta q^{\theta}(s_{k},a_{k})-\log\sum_{b}\exp\left(\beta q^{\theta}(s_{k},b)\right)\right).\end{split}$
Now, denote a perturbed version of the $q^{\theta}$-parameterized online dataset by
~$\tilde{\mathcal{H}}_{t}(q^{\theta})=\{(s_{k},a_{k},\check{s}_{k},\bar{r}_{k},\tilde{\mathcal{E}}_{k})_{k=N:N+t}\},$
and thus similar to before, we have
$\begin{split}&\quad~{}\log P(\tilde{\mathcal{H}}_{t}(q^{\theta})|\tilde{\mathcal{D}}_{0}(q^{\theta}),q^{\theta})\\ &=\sum_{k=N}^{N+t}\left(\log P(\tilde{\mathcal{E}}_{k}(q^{\theta})|\check{s}_{k},a_{k},s_{k},\bar{r}_{k},q^{\theta})+\log P(\check{s}_{k}|a_{k},s_{k},q^{\theta})\right),\\ &\leq\sum_{k=N}^{N+t}\log P(\tilde{\mathcal{E}}_{k}|\check{s}_{k},a_{k},s_{k},\bar{r}_{k},q^{\theta}),\\ \end{split}$
where we again ignored the transition term to obtain an optimistic upper bound.
Given the random perturbations above, we have
$\begin{split}&\quad~{}\log P(\tilde{\mathcal{E}}_{k}(q^{\theta})|\check{s}_{k},a_{k},s_{k},\bar{r}_{k},q^{\theta})\\ &=-\frac{1}{2}\left(r_{k}+z_{k}+\max_{b}q^{\theta}(\check{s}_{k},b)-q^{\theta}(s_{k},a_{k})-\bar{r}_{k}\right)^{2}+\text{constant}.\end{split}$
The prior over $\beta$, $f_{2}(\beta)$ is assumed to be an exponential pdf $\lambda_{2}\exp(-\lambda_{2}\beta),\beta\geq 0$, while that over $\theta$ is assumed Gaussian. Thus, putting it all together, we get the following optimistic loss function (to minimize over $\theta$ and $\beta$),
~$\begin{split}\tilde{\mathcal{L}}(\theta,\beta)&=\frac{1}{2\sigma^{2}}\sum_{k=0}^{N+T-1}\left(r_{k}+z_{k}+\max_{b}q^{\theta}(\check{s}_{k},b)-q^{\theta}(s_{k},a_{k})-\bar{r}_{k}\right)^{2}\\ &-\sum_{t=0}^{T-1}w_{k}\left(\beta q^{\theta}(s_{t},a_{t})-\log\sum_{b}\exp\left(\beta q^{\theta}(s_{t},b)\right)\right)+\frac{1}{2}(\theta-\tilde{\theta})^{\top}\Sigma_{0}(\theta-\tilde{\theta})+\lambda_{2}\beta.\end{split}$ (22)
The above loss function is difficult to optimize in general due to the $\max$ operation, and the $Q$-value function in general having a nonlinear form.
Remark 4.1.
Note that the loss function in (22) can be hard to jointly optimize over $\theta$ and $\beta$. In particular, estimates of $\beta$ can be quite noisy when $\beta$ is large, and the near-optimal expert policy only covers the state-action space partially. Thus, we consider other methods of estimating $\beta$ that are more robust, which can then be plugged into the loss function in (22). Specifically, we could simply look at the entropy of the empirical distribution of the actions in the offline dataset. Suppose the empirical distribution of $\{\bar{a}_{0}^{l},\ldots\bar{a}_{k}\}_{l=1}^{L}$ is $\mu_{A}$. Then we use $c_{0}/\mathcal{H}(\mu_{A})$ as an estimation for $\beta$, where $c_{0}>0$ is a hyperparameter. The intuition is that for smaller $\beta$, the offline actions tend to be more uniform and thus the entropy will be large. This is an unsupervised approach and agnostic to specific offline data generation process.
Remark 4.2.
In the loss function in (22), the parameter $\theta$ appears inside the $\max$ operation. Thus, it can be quite difficult to optimize over $\beta$. Since the loss function is typically optimized via an iterative algorithm such as a gradient descent method, a simple and scalable solution that works well in practice is to use the parameter estimate $\theta$ from the previous iteration inside the $\max$ operation, and thus optimize over $\theta$ only in the other terms.
4.2 inf-iRLSVI  bridges Online RL and Imitation Learning
In the previous subsection, we derived inf-iRLSVI, a Bayesian-bootstrapped algorithm. We now present interpretation of the algorithm as bridging online RL (via commonality with the RLSVI algorithm (Osband et al., 2016) and imitation learning, and hence a way for its generalization.
Consider the RLSVI algorithm for online reinforcement learning as introduced in Osband et al. (2019). It draws its inspiration from the posterior sampling principle for online learning, and has excellent cumulative regret performance. The RLSVI algorithm, that uses all of the data available at the end of episode $t$, including any offline dataset involves minimizing the corresponding loss function at each time step:
~$\begin{split}\tilde{\mathcal{L}}_{\text{RLSVI}}(\theta)&=\frac{1}{2\sigma^{2}}\sum_{k=1}^{L+t}\sum_{h=0}^{H-1}\left(r_{k}+\max_{b}Q^{\theta}_{t+1}(\check{s}_{k},b)-q^{\theta}_{t}(s_{k},a_{k})\right)^{2}\\ &+\frac{1}{2}({\theta_{0:H}}-{\tilde{\theta}_{0:H}})^{\top}\Sigma_{0}({\theta_{0:H}}-{\tilde{\theta}_{0:H}}).\end{split}$
Now, let us consider an imitation learning setting. Let $\tau_{l}=(s_{k},a_{k},\check{s}_{k})_{h=0}^{H-1}$ be the trajectory of the $l$th episode. Let $\hat{\pi}_{t}(a|s)$ denote the empirical estimate of probability of taking action $a$ in state $s$ at time $h$, i.e., an empirical estimate of the expert’s randomized policy. Let $p(\tau)$ denote the probability of observing the trajectory under the policy $\hat{\pi}$.
Let $\pi_{t}^{\beta,\theta}(\cdot|s)$ denote the parametric representation of the policy used by the expert. And let $p^{\beta,\theta}(\tau)$ denote the probability of observing the trajectory $\tau$ under the policy ${\pi}^{\beta,\theta}$. Then, the loss function corresponding to the KL divergence between $\Pi_{l=1}^{L}p(\tau_{l})$ and $\Pi_{l=1}^{L}p^{\beta,\theta}(\tau_{l})$ is given by
~$\tilde{\mathcal{L}}_{\text{IL}}(\beta,\theta)$ $=D_{KL}\left(\Pi_{l=1}^{L}p(\tau_{l})||\Pi_{l=1}^{L}p^{\beta,\theta}(\tau_{l})\right)=\int\Pi_{l=1}^{L}p(\tau_{l})\log\frac{\Pi_{l=1}^{L}p(\tau_{l})}{\Pi_{l=1}^{L}p^{\beta}(\tau_{l})}$ (23)
$=\sum_{l=1}^{L}\int p(\tau_{l})\log\frac{p(\tau_{l})}{p^{\beta,\theta}(\tau_{l})},=\sum_{l=1}^{L}\sum_{h=0}^{H-1}\log\frac{\hat{\pi}_{t}(a_{k}|s_{k})}{\pi^{\beta,\theta}_{t}(a_{k}|s_{k})}$ (24)
$=\sum_{l=1}^{L}\sum_{h=0}^{H-1}[\log\hat{\pi}_{t}(a_{k}|s_{k})-\log\pi_{t}^{\beta,\theta}(a_{k}|s_{k})]$ (25)
$=-\sum_{l=1}^{L}\sum_{h=0}^{H-1}\left(\beta q^{\theta}_{t}(s_{k},a_{k})-\log\sum_{b}\exp\left(\beta q^{\theta}_{t}(s_{k},b)\right)\right)+\text{constant}.$ (26)
Remark 4.3.
(i) The loss function ~$\tilde{\mathcal{L}}_{\text{IL}}(\beta,\theta)$ is the same as the second (action-likelihood) term in (22) while the loss function ~$\tilde{\mathcal{L}}_{\text{RLSVI}}(\theta)$ is the same as the first and third terms there (except for perturbation) and minus the $\lambda_{2}\beta$ term that corresponds to the prior over $\beta$. (ii) Note that while we used the more common KL divergence for the imitation learning loss function, use of log loss would yield the same outcome.
Thus, the inf-iRLSVI loss function can be viewed as
~$\tilde{\mathcal{L}}(\beta,\theta)=\tilde{\mathcal{L}}_{\text{RLSVI}}(\theta)+\tilde{\mathcal{L}}_{\text{IL}}(\beta,\theta)+\lambda_{2}\beta,$ (27)
thus establishing that the proposed algorithm may be viewed as bridging Online RL with Imitation Learning. Note that the last term corresponds to the prior over $\beta$. If $\beta$ is known (or uniform), it will not show up in the loss function above.
The above also suggests a possible way to generalize and obtain other online learning algorithms that can bootstrap by use of offline datasets. Namely, at each step, they can optimize a general loss function of the following kind:
~$\tilde{\mathcal{L}}_{\alpha}(\beta,\theta)=\alpha\tilde{\mathcal{L}}_{\text{ORL}}(\theta)+(1-\alpha)\tilde{\mathcal{L}}_{\text{IL}}(\beta,\theta)+\lambda_{2}\beta,$ (28)
where ~$\tilde{\mathcal{L}}_{\text{ORL}}$ is a loss function for an Online RL algorithm, ~$\tilde{\mathcal{L}}_{\text{IL}}$ is a loss function for some Imitation Learning algorithm, and factor $\alpha\in[0,1]$ provides a way to tune between emphasizing the offline imitation learning and the online reinforcement learning.
5 Conclusions
In this paper, we have introduced an ideal Bayesian online reinforcement learning algorithm, inf-iPSRL, that naturally uses any offline data available for pre-training to boot-strap online learning. The algorithm is agnostic to whether we have a small or large amount of data, or a small or large window for online learning. It is also useful in the special case of no offline data but not for the ‘pure’ offline RL setting when there is no online learning phase. A key aspect of the proposed algorithm is that unlike past PSRL algorithms, our algorithm takes the episode length as an input. Our analysis then shows what type of episodic schedules result in near-optimal regret. The analysis involves several novel elements and depends on prior-dependent regret analysis, which is known to be quite challenging and itself a novelty.
We then presented an approximate Bayesian online learning algorithm, inf-iRLSVI that can be seen as naturally bridging online reinforcement learning and imitation learning. In prior work (Osband et al., 2019), such algorithms are well-known to be computationally practical in problems that need deep exploration.
We note that while it should be possible to design non-Bayesian algorithms for the problem we address in this paper, and analyze its frequentist regret, it is much less natural to incorporate information derived from offline datasets in such algorithms, and hence our choice to focus on Bayesian algorithms with a guarantee on their Bayesian regret.
A number of future directions are possible: The performance of the inf-iRLSVI algorithm can be empirically and theoretically be analyzed. Theoretical analysis is expected to be a combination of analysis of RLSVI in Osband et al. (2019) along with novel elements we have introduced in the analysis of the inf-iPSRL algorithm in this paper. This is left as a useful exercise for aspiring student researchers. Yet another future direction is design of universal RL algorithms which can specialize to Online RL, or be seen as Offline RL or Imitation Learning algorithms in suitable settings, or anything in between. Currently, we have to design RL algorithms separately for the three settings.
Appendix A Proof of Theorem 3.1
Recall that $t_{k}=\sum_{l=1}^{k-1}T_{k}$ is the starting time of learning episode $k$. Let $\tilde{\mathcal{D}}_{k}:=\mathcal{D}_{0}\cup\{(s_{0},a_{0},\cdots,s_{t_{k}-1},a_{t_{k}-1},s_{t_{k}})\}$ denote the offline and online data the learning agent used to form a posterior to sample the environment for the $k$-th learning episode. For simplicity of notation, let $\mathbb{P}_{k}(\cdot)=\mathbb{P}(\cdot|\tilde{\mathcal{D}}_{k})$ and $\mathbb{E}_{k}[\cdot]=\mathbb{E}[\cdot|\tilde{\mathcal{D}}_{k}]$. Also define $\tilde{\mathbb{E}}_{k}^{*}[\cdot]=\mathbb{E}[\cdot|\tilde{\mathcal{D}}_{k},\theta^{*},\tilde{\theta}^{k}]$.
For convenience, we fix $T$ and write $K=K_{T}$. For the sake of this proof, if the last episode goes beyond time $T-1$, we “cut off” the episode after time $T-1$ and define $T_{K}=T-1-t_{K}$ and $t_{K+1}=T$.
Without loss of generality, by choosing an appropriate reference state $\hat{s}$ in (3) we can assume that $v(s;\theta)\in[0,\bar{v}]$.
We will start with the following lemma which bounds all the mismatch probabilities by the initial mismatch probability.
Lemma A.1.
$\mathbb{P}(\tilde{\pi}^{k}\neq\pi^{*})\leq\mathbb{P}(\tilde{\pi}^{1}\neq\pi^{*})$ for all $k\geq 1$.
Proof.
Define $f(x)=x(1-x)$. We have
$\mathbb{P}(\tilde{\pi}^{k}\neq\pi^{*})$ $=\mathbb{E}[\mathbb{P}_{k}(\tilde{\pi}^{k}\neq\pi^{*})]=\mathbb{E}\left[\sum_{\pi\in\varPi}\mathbb{P}_{k}(\tilde{\pi}^{k}=\pi,\pi^{*}\neq\pi)\right]$ (29)
$=\mathbb{E}\left[\sum_{\pi\in\varPi}f(\mathbb{P}_{k}(\pi^{*}=\pi))\right]$ (30)
$=\mathbb{E}\left[\sum_{\pi\in\varPi}\mathbb{E}_{1}[f(\mathbb{P}_{k}(\pi^{*}=\pi))]\right]\leq\mathbb{E}\left[\sum_{\pi\in\varPi}f(\mathbb{E}_{1}[\mathbb{P}_{k}(\pi^{*}=\pi)])\right]$ (31)
$=\mathbb{E}\left[\sum_{\pi\in\varPi}f(\mathbb{P}_{1}(\pi^{*}=\pi))\right]=\mathbb{E}\left[\sum_{\pi\in\varPi}\mathbb{P}_{1}(\tilde{\pi}^{1}=\pi,\pi^{*}\neq\pi)\right]$ (32)
$=\mathbb{P}(\tilde{\pi}^{1}\neq\pi^{*})$ (33)
where the inequality is due to Jensen’s Inequality and the fact that $f$ is a concave function. ∎
Define the episode $k$ regret $Z_{k}$ as
$Z_{k}=T_{k}\bar{r}^{*}(\theta^{*})-\sum_{t=t_{k}}^{t_{k+1}-1}r(s_{t},a_{t}).$ (34)
Then we can write $\mathfrak{BR}_{T}(\phi^{\mathrm{iPSRL}})=\sum_{k=1}^{K}\mathbb{E}[Z_{k}]$.
If the event $\{\tilde{\pi}^{k}=\pi^{*}\}$ is true, then for $t_{k}\leq t<t_{k+1}$, we have $a_{t}=\pi^{*}(s_{t})$ and $v(s_{t};\theta^{*})+\bar{r}^{*}(\theta^{*})=r(s_{t},a_{t})+\sum_{s^{\prime}\in\mathcal{S}}\theta^{*}(s^{\prime}|s_{t},a_{t})v(s^{\prime};\theta^{*})$ due to the assumption on the planning oracle MDPSolve. Therefore,
$Z_{k}\bm{1}_{\{\tilde{\pi}^{k}=\pi^{*}\}}=\sum_{t=t_{k}}^{t_{k+1}-1}\left(\sum_{s^{\prime}\in\mathcal{S}}\theta^{*}(s^{\prime}|s_{t},a_{t})v(s^{\prime};\theta^{*})-v(s_{t};\theta^{*})\right)\bm{1}_{\{\tilde{\pi}^{k}=\pi^{*}\}}$ (35)
Furthermore, using the fact that $\tilde{\mathbb{E}}_{k}^{*}[v(s_{t+1};\theta^{*})|s_{t},a_{t}]=\sum_{s^{\prime}\in\mathcal{S}}\theta^{*}(s^{\prime}|s_{t},a_{t})v(s^{\prime};\theta^{*})$, we have
$\quad~{}\tilde{\mathbb{E}}_{k}^{*}[Z_{k}\bm{1}_{\{\tilde{\pi}^{k}=\pi^{*}\}}]$ (36)
$=\tilde{\mathbb{E}}_{k}^{*}\left[\sum_{t=t_{k}}^{t_{k+1}-1}\left(\sum_{s^{\prime}\in\mathcal{S}}\theta^{*}(s^{\prime}|s_{t},a_{t})v(s^{\prime};\theta^{*})-v(s_{t};\theta^{*})\right)\bm{1}_{\{\tilde{\pi}^{k}=\pi^{*}\}}\right]$ (37)
$=\tilde{\mathbb{E}}_{k}^{*}\left[\sum_{t=t_{k}}^{t_{k+1}-1}\left(\sum_{s^{\prime}\in\mathcal{S}}\theta^{*}(s^{\prime}|s_{t},a_{t})v(s^{\prime};\theta^{*})-v(s_{t};\theta^{*})\right)\right]\bm{1}_{\{\tilde{\pi}^{k}=\pi^{*}\}}$ (38)
$=\tilde{\mathbb{E}}_{k}^{*}\left[\sum_{t=t_{k}}^{t_{k+1}-1}\left(v(s_{t+1};\theta^{*})-v(s_{t};\theta^{*})\right)\right]\bm{1}_{\{\tilde{\pi}^{k}=\pi^{*}\}}$ (39)
$=\tilde{\mathbb{E}}_{k}^{*}\left[v(s_{t_{k+1}};\theta^{*})-v(s_{t_{k}};\theta^{*})\right]\bm{1}_{\{\tilde{\pi}^{k}=\pi^{*}\}}$ (40)
$=\tilde{\mathbb{E}}_{k}^{*}\left[\left[v(s_{t_{k+1}};\theta^{*})-v(s_{t_{k}};\theta^{*})\right]\bm{1}_{\{\tilde{\pi}^{k}=\pi^{*}\}}\right]$ (41)
Taking the sum over all episodes $k$, we have
$\quad~{}\sum_{k=1}^{K}\mathbb{E}\left[Z_{k}\bm{1}_{\{\tilde{\pi}^{k}=\pi^{*}\}}\right]$ (42)
$=\mathbb{E}\left[\sum_{k=1}^{K}[v(s_{t_{k+1}};\theta^{*})-v(s_{t_{k}};\theta^{*})]\bm{1}_{\{\tilde{\pi}^{k}=\pi^{*}\}}\right]$ (43)
$=\mathbb{E}\left[v(s_{t_{K+1}};\theta^{*})\bm{1}_{\{\tilde{\pi}^{K}=\pi^{*}\}}-v(s_{t_{1}};\theta^{*})\bm{1}_{\{\tilde{\pi}^{1}=\pi^{*}\}}\right]$ (44)
$\quad+\mathbb{E}\left[\sum_{k=1}^{K-1}v(s_{t_{k+1}};\theta^{*})\bm{1}_{\{\tilde{\pi}^{k}=\pi^{*},\tilde{\pi}^{k+1}\neq\pi^{*}\}}-\sum_{k=2}^{K}v(s_{t_{k}};\theta^{*})\bm{1}_{\{\tilde{\pi}^{k-1}\neq\pi^{*},\tilde{\pi}^{k}=\pi^{*}\}}\right]$ (45)
$\leq\bar{v}+\bar{v}\mathbb{E}\left[\sum_{k=1}^{K-1}\bm{1}_{\{\tilde{\pi}^{k}=\pi^{*},\tilde{\pi}^{k+1}\neq\pi^{*}\}}\right]\leq\bar{v}+\bar{v}\sum_{k=1}^{K-1}\mathbb{P}(\tilde{\pi}^{k+1}\neq\pi^{*})$ (46)
$\leq\bar{v}+\bar{v}(K-1)\varepsilon$ (47)
where we used Lemma A.1 in the last inequality.
In general, for $t_{k}\leq t<t_{k+1}$, since $a_{t}=\tilde{\pi}^{k}(s_{t})$, we have $v(s_{t};\tilde{\theta}^{k})+\bar{r}^{*}(\tilde{\theta}^{k})=r(s_{t},a_{t})+\sum_{s^{\prime}\in\mathcal{S}}\tilde{\theta}^{k}(s^{\prime}|s_{t},a_{t})v(s^{\prime};\tilde{\theta}^{k})$ due to the assumption on the planning oracle MDPSolve. Therefore,
$Z_{k}$ $=T_{k}\bar{r}^{*}(\theta^{*})-\sum_{t=t_{k}}^{t_{k+1}-1}\left(\bar{r}^{*}(\tilde{\theta}^{k})+v(s_{t};\tilde{\theta}^{k})-\sum_{s^{\prime}\in\mathcal{S}}\tilde{\theta}^{k}(s^{\prime}|s_{t},a_{t})v(s^{\prime};\tilde{\theta}^{k})\right)$ (48)
$=\underbrace{T_{k}\left[\bar{r}^{*}(\theta^{*})-\bar{r}^{*}(\tilde{\theta}^{k})\right]}_{Z_{k,0}}+\underbrace{\sum_{t=t_{k}}^{t_{k+1}-1}\left(\sum_{s^{\prime}\in\mathcal{S}}\theta^{*}(s^{\prime}|s_{t},a_{t})v(s^{\prime};\tilde{\theta}^{k})-v(s_{t};\tilde{\theta}^{k})\right)}_{Z_{k,1}}$ (49)
$\quad+\underbrace{\sum_{t=t_{k}}^{t_{k+1}-1}\sum_{s^{\prime}\in\mathcal{S}}[\tilde{\theta}^{k}(s^{\prime}|s_{t},a_{t})-\theta^{*}(s^{\prime}|s_{t},a_{t})]v(s^{\prime};\tilde{\theta}^{k})}_{Z_{k,2}}.$ (50)
We will analyze the three terms in (50) separately. First, we have
$\mathbb{E}_{k}[Z_{k,0}\bm{1}_{\{\tilde{\pi}^{k}\neq\pi^{*}\}}]=T_{k}\left(\mathbb{E}_{k}[\bar{r}^{*}(\theta^{*})\bm{1}_{\{\tilde{\pi}^{k}\neq\pi^{*}\}}]-\mathbb{E}_{k}[\bar{r}^{*}(\tilde{\theta}^{k})\bm{1}_{\{\pi^{*}\neq\tilde{\pi}^{k}\}}]\right)=0$ (51)
since $\theta^{*}$ and $\tilde{\theta}^{k}$ are i.i.d. given $\tilde{\mathcal{D}}_{k}$. Following similar steps as (41) we have
$\tilde{\mathbb{E}}_{k}^{*}[Z_{k,1}\bm{1}_{\{\tilde{\pi}^{k}\neq\pi^{*}\}}]=\tilde{\mathbb{E}}_{k}^{*}\left[v(s_{t_{k+1}};\tilde{\theta}^{k})-v(s_{t_{k}};\tilde{\theta}^{k})\right]\bm{1}_{\{\tilde{\pi}^{k}\neq\pi^{*}\}}\leq\bar{v}\bm{1}_{\{\tilde{\pi}^{k}\neq\pi^{*}\}}$ (52)
and hence
$\mathbb{E}[Z_{k,1}\bm{1}_{\{\tilde{\pi}^{k}\neq\pi^{*}\}}]\leq\bar{v}\mathbb{P}(\tilde{\pi}^{k}\neq\pi^{*})\leq\bar{v}\varepsilon.$ (53)
where we used Lemma A.1. Combining (47)(50)(51)(53), we have
$\sum_{k=1}^{K}\mathbb{E}[Z_{k}]$ $=\sum_{k=1}^{K}\mathbb{E}[Z_{k}\bm{1}_{\{\tilde{\pi}^{k}=\pi^{*}\}}]+\sum_{k=1}^{K}\mathbb{E}[(Z_{k,0}+Z_{k,1}+Z_{k,2})\bm{1}_{\{\tilde{\pi}^{k}\neq\pi^{*}\}}]$ (54)
$\leq\bar{v}+\bar{v}(K-1)\varepsilon+\sum_{k=1}^{K}\mathbb{E}[Z_{k,1}\bm{1}_{\{\tilde{\pi}^{k}\neq\pi^{*}\}}]+\sum_{k=1}^{K}\mathbb{E}[Z_{k,2}\bm{1}_{\{\tilde{\pi}^{k}\neq\pi^{*}\}}]$ (55)
$\leq\bar{v}+\bar{v}(2K-1)\varepsilon+\sum_{k=1}^{K}\mathbb{E}[Z_{k,2}\bm{1}_{\{\tilde{\pi}^{k}\neq\pi^{*}\}}]$ (56)
Now we bound the summation term in (56). For convenience, define $I_{k}=\bm{1}_{\{\tilde{\pi}^{k}\neq\pi^{*}\}}$. First, since $v(s;\theta)\in[0,\bar{v}]$, we have
$Z_{k,2}$ $\leq\bar{v}\sum_{t=t_{k}}^{t_{k+1}-1}\sum_{s^{\prime}\in\mathcal{S}}(\tilde{\theta}^{k}(s^{\prime}|s_{t},a_{t})-\theta^{*}(s^{\prime}|s_{t},a_{t}))_{+}$ (57)
$=\bar{v}\sum_{t=t_{k}}^{t_{k+1}-1}\frac{1}{2}\|\tilde{\theta}^{k}(\cdot|s_{t},a_{t})-\theta^{*}(\cdot|s_{t},a_{t})\|_{1}$ (58)
Following Osband et al. (2013); Ouyang et al. (2017), define $N_{t}(s,a):=\sum_{\tau=0}^{t-1}\bm{1}_{\{(s_{\tau},a_{\tau})=(s,a)\}}$ to be the number of occurrences of the state-action pair $(s,a)$ from time 00 to time $t-1$.
Claim: For any $k,t$ such that $t_{k}\leq t<t_{k+1}$ and any $\delta\in(0,1)$,
$\quad~{}\mathbb{E}\left[\frac{1}{2}\|\tilde{\theta}^{k}(\cdot|s_{t},a_{t})-\theta^{*}(\cdot|s_{t},a_{t})\|_{1}\cdot I_{k}\right]$ (59)
$\leq\mathbb{E}\left[\min\left\{\sqrt{\dfrac{(2\log 2)S+2\log(1/\delta)}{N_{t_{k}}(s_{t},a_{t})}},2\right\}\cdot I_{k}\right]+2t_{k}SA\delta$ (60)
For $t_{k}\leq t<t_{k+1}$, define the event
$\mathcal{E}_{t}=\{N_{t}(s_{t},a_{t})>2N_{t_{k}}(s_{t},a_{t})\}$ (61)
We can decompose the first term in (60) by
$\quad~{}\min\left\{\sqrt{\dfrac{(2\log 2)S+2\log(1/\delta)}{N_{t_{k}}(s_{t},a_{t})}},2\right\}I_{k}$ (62)
$\leq\min\left\{\sqrt{\dfrac{(2\log 2)S+2\log(1/\delta)}{N_{t_{k}}(s_{t},a_{t})}},2\right\}I_{k}\bm{1}_{\mathcal{E}_{t}^{c}}+2I_{k}\bm{1}_{\mathcal{E}_{t}}$ (63)
$\leq\min\left\{\sqrt{\dfrac{(4\log 2)S+4\log(1/\delta)}{N_{t}(s_{t},a_{t})}},2\right\}I_{k}\bm{1}_{\mathcal{E}_{t}^{c}}+2I_{k}\bm{1}_{\mathcal{E}_{t}}$ (64)
$\leq\underbrace{\sqrt{\dfrac{(4\log 2)S+4\log(1/\delta)}{\max\{N_{t}(s_{t},a_{t}),1\}}}}_{\beta_{t}}I_{k}+2I_{k}\bm{1}_{\mathcal{E}_{t}}$ (65)
Combining (58)(60)(65) and set $\delta=(SAKT)^{-1}$, we have
$\quad~{}\sum_{k=1}^{K}\mathbb{E}[Z_{k,2}I_{k}]$ (66)
$\leq\bar{v}\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=t_{k}}^{t_{k+1}-1}\beta_{t}I_{k}\right]+2\bar{v}\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=t_{k}}^{t_{k+1}-1}I_{k}\bm{1}_{\mathcal{E}_{t}}\right]+2\bar{v}SA\delta\sum_{k=1}^{K}t_{k}$ (67)
$\leq\bar{v}\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=t_{k}}^{t_{k+1}-1}\beta_{t}I_{k}\right]+2\bar{v}\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=t_{k}}^{t_{k+1}-1}I_{k}\bm{1}_{\mathcal{E}_{t}}\right]+2\bar{v}$ (68)
By Cauchy-Schwatz inequality, we have
$\quad~{}\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=t_{k}}^{t_{k+1}-1}\beta_{t}I_{k}\right]$ (69)
$\leq\sqrt{\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=t_{k}}^{t_{k+1}-1}I_{k}^{2}\right]}\sqrt{\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=t_{k}}^{t_{k+1}-1}\beta_{t}^{2}\right]}$ (70)
$=\sqrt{\sum_{k=1}^{K}\sum_{t=t_{k}}^{t_{k+1}-1}\mathbb{E}[I_{k}]}\sqrt{\mathbb{E}\left[\sum_{t=0}^{T-1}\beta_{t}^{2}\right]}\stackrel{{\scriptstyle(a)}}{{\leq}}\sqrt{\varepsilon T}\sqrt{\mathbb{E}\left[\sum_{t=0}^{T-1}\beta_{t}^{2}\right]}$ (71)
$=\sqrt{\varepsilon T[(4\log 2)S+4\log(SAKT)]\mathbb{E}\left[\sum_{t=0}^{T-1}\left(\max\{N_{t}(s_{t},a_{t}),1\}\right)^{-1}\right]}$ (72)
where in $(a)$ we have used Lemma A.1.
Furthermore, we have
$\quad~{}\sum_{t=0}^{T-1}\left(\max\{N_{t}(s_{t},a_{t}),1\}\right)^{-1}=\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\sum_{j=0}^{N_{T}(s,a)-1}\left(\max\{j,1\}\right)^{-1}$ (73)
$\leq\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left[1+\log(N_{T}(s,a)+1)\right]\leq SA\left[1+\log\left(\frac{T}{SA}+1\right)\right]$ (74)
where the last inequality is due to the concavity of the function $f(x)=\log(x+1)$.
Fix a state-action pair $(s,a)\in\mathcal{S}\times\mathcal{A}$, let $k_{1}<k_{2}<...<k_{m}$ be the first $m$ episodes such that the event $\{N_{t}(s,a)>2N_{t_{k}}(s,a)\}$ occurs. Let $\bar{\tau}_{l}:=t_{k_{l}+1}-1$ denote the last timestamp in episode $k_{l}$. Using induction, it can be shown that $N_{\bar{\tau}_{l}}(s,a)\geq 2^{l}-1$ for $1\leq l\leq m$. Therefore, the above event can occur in no more than $\log_{2}(N_{T}(s,a)+1)$ episodes. Therefore, we have
$\sum_{k=1}^{K}\sum_{t=t_{k}}^{t_{k+1}-1}\bm{1}_{\mathcal{E}_{t}}$ $\leq\sum_{(s,a)\in\mathcal{S}\times\mathcal{A}}\left(\max_{k\in[K]}T_{k}\right)\log_{2}(N_{T}(s,a)+1)$ (75)
$\leq SA\left(\max_{k\in[K]}T_{k}\right)\log_{2}\left(\frac{T}{SA}+1\right).$ (76)
where the last inequality is due to the concavity of the function $f(x)=\log(x+1)$.
Also note that one can obtain a prior-dependent bound on the second term in (68) can be bounded with Lemma A.1:
$\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=t_{k}}^{t_{k+1}-1}I_{k}\bm{1}_{\mathcal{E}_{t}}\right]\leq\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=t_{k}}^{t_{k+1}-1}I_{k}\right]=\sum_{k=1}^{K}\sum_{t=t_{k}}^{t_{k+1}-1}\mathbb{E}\left[I_{k}\right]\leq\varepsilon T$ (77)
Therefore, we conclude that
$\mathbb{E}\left[\sum_{k=1}^{K}\sum_{t=t_{k}}^{t_{k+1}-1}I_{k}\bm{1}_{\mathcal{E}_{t}}\right]\leq\min\left\{\varepsilon T,SA\left(\max_{k\in[K]}T_{k}\right)\log_{2}\left(\frac{T}{SA}+1\right)\right\}$ (78)
Finally, combining (56)(68)(72)(74)(78), we have
$\quad~{}\sum_{k=1}^{K}\mathbb{E}[Z_{k}]$ (79)
$\leq\bar{v}+\bar{v}\varepsilon(2K-1)+\bar{v}\sqrt{\varepsilon T[(4\log 2)S+4\log(SAKT)]SA\left[1+\log\left(\frac{T}{SA}+1\right)\right]}+$ (80)
$+2\bar{v}\min\left\{\varepsilon T,SA\left(\max_{k\in[K]}T_{k}\right)\log_{2}\left(\frac{T}{SA}+1\right)\right\}+2\bar{v}$ (81)
$\leq 3\bar{v}+2\bar{v}(R_{1}+R_{2}+R_{3})$ (82)
where
$R_{1}$ $:=\varepsilon K$ (83)
$R_{2}$ $:=\sqrt{\varepsilon S^{2}AT\log(2SAKT)\left[1+\log\left(\frac{T}{SA}+1\right)\right]}$ (84)
$R_{3}$ $:=\min\left\{\varepsilon T,SA\left(\max_{k\in[K]}T_{k}\right)\log_{2}\left(\frac{T}{SA}+1\right)\right\}$ (85)
Proof of Claim.
Let $\check{\theta}^{n}(\cdot|s,a)$ be the empirical distribution of the next states of the state-action pair $(s,a)$ after observing $(s,a)$ exactly $n$ times. By the $L_{1}$ concentration inequality for empirical distributions (Weissman et al., 2003), we have
$\mathbb{P}_{\theta^{*}}(\|\check{\theta}^{n}(\cdot|s,a)-\theta^{*}(\cdot|s,a)\|_{1}>\lambda)\leq 2^{S}\exp\left(-\frac{1}{2}n\lambda^{2}\right)$ (86)
for any $\lambda>0$. For $\delta>0$, define $\xi(n,\delta):=\min\left\{\sqrt{\frac{(2\log 2)S+2\log(1/\delta)}{n}},2\right\}$, we have
$\mathbb{P}_{\theta^{*}}\left(\|\check{\theta}^{n}(\cdot|s,a)-\theta^{*}(\cdot|s,a)\|_{1}>\xi(n,\delta)\right)\leq\delta.$ (87)
Define
$\hat{\theta}^{t}(\cdot|s,a):=\begin{cases}\frac{1}{t}\sum_{\tau=0}^{t-1}\bm{1}_{\{(s_{\tau},a_{\tau})=(s,a),s_{\tau+1}=\cdot\}}&t>0\\ \text{uniform}&t=0\end{cases}$ (88)
to be the empirical distribution of next states of $(s,a)$ in the first $t$ steps. Define
$\hat{\varTheta}^{k}$ =$=\left\{\theta:\|\hat{\theta}^{t_{k}}(\cdot|s,a)-\theta(\cdot|s,a)\|_{1}\leq\xi(N_{t_{k}}(s,a),\delta)~{}~{}\forall s\in\mathcal{S},a\in\mathcal{A}\right\}$
Then, we have $\mathbb{P}(\theta^{*}\not\in\hat{\varTheta}^{k})=0$ if $t_{k}=0$ (since $\xi(0,\delta)=2$) and
$\quad~{}\mathbb{P}(\theta^{*}\not\in\hat{\varTheta}^{k})$ (89)
$\leq\mathbb{P}\left(\exists n\in[t_{k}],s\in\mathcal{S},a\in\mathcal{A}~{}~{}\|\check{\theta}^{n}(\cdot|s,a)-\theta^{*}(\cdot|s,a)\|_{1}>\xi(n,\delta)\right)$ (90)
$\leq t_{k}SA\delta$ (91)
if $t_{k}>0$.
Since $\theta^{*}$ and $\tilde{\theta}^{k}$ are i.i.d. given $\tilde{\mathcal{D}}_{k}$, and the random set $\hat{\varTheta}^{k}$ is measurable with respect to $\tilde{\mathcal{D}}_{k}$, we have $\mathbb{P}(\theta^{*}\not\in\hat{\varTheta}^{k})=\mathbb{P}(\tilde{\theta}^{k}\not\in\hat{\varTheta}^{k})$. Therefore we have
$\quad~{}\mathbb{E}\left[\|\tilde{\theta}^{k}(\cdot|s_{t},a_{t})-\theta^{*}(\cdot|s_{t},a_{t})\|_{1}\cdot I_{k}\right]$ (92)
$\leq\mathbb{E}\left[\|\tilde{\theta}^{k}(\cdot|s_{t},a_{t})-\theta^{*}(\cdot|s_{t},a_{t})\|_{1}\cdot I_{k}\bm{1}_{\{\theta^{*},\tilde{\theta}^{k}\in\hat{\varTheta}^{k}\}}\right]+$ (93)
$+\mathbb{E}\left[\|\tilde{\theta}^{k}(\cdot|s_{t},a_{t})-\theta^{*}(\cdot|s_{t},a_{t})\|_{1}\cdot I_{k}\left(\bm{1}_{\{\theta^{*}\not\in\hat{\varTheta}^{k}\}}+\bm{1}_{\{\tilde{\theta}^{k}\not\in\hat{\varTheta}^{k}\}}\right)\right]$ (94)
$\leq\mathbb{E}[2\xi(N_{t_{k}}(s_{t},a_{t}),\delta)\cdot I_{k}]+2\left(\mathbb{P}(\theta^{*}\not\in\hat{\varTheta}^{k})+\mathbb{P}(\tilde{\theta}^{k}\not\in\hat{\varTheta}^{k})\right)$ (95)
$\leq 2\mathbb{E}[\xi(N_{t_{k}}(s_{t},a_{t}),\delta)\cdot I_{k}]+4t_{k}SA\delta$ (96)
Appendix B Proof of Lemma 3.8
In this proof, let $\mathbb{P}_{\theta^{*}}$ be the distribution induced by the expert’s policy $\pi^{\beta}$ conditioning on $\theta^{*}$.
Under Assumption 2.2, we have $q(s,a^{\prime};\theta)-q(s,a;\theta)\leq r(s,a^{\prime})-r(s,a)+\bar{v}\leq 1+\bar{v}$ for any $s\in\mathcal{S},a,a^{\prime}\in\mathcal{A}$. Therefore, for any $s\in\mathcal{S},a\in\mathcal{A}$, we have
$\pi^{\beta}(a|s)$ $:=\dfrac{\exp(\beta q(s,a;\theta^{*}))}{\sum_{a^{\prime}\in\mathcal{A}}\exp(\beta q(s,a^{\prime};\theta^{*}))}$ (97)
$=\dfrac{1}{\sum_{a^{\prime}\in\mathcal{A}}\exp[\beta(q(s,a^{\prime};\theta^{*})-q(s,a;\theta^{*}))]}$ (98)
$\geq\dfrac{1}{1+(A-1)\exp(\beta(1+\bar{v}))}=:\pi_{\min}$ (99)
Combining (99) with Assumption 3.7, for any $s,s^{\prime}\in\mathcal{S}$,
$\mathbb{P}_{\theta^{*}}(\exists 0<k\leq S,s_{k}=s^{\prime}|s_{0}=s)\geq\delta(\pi_{\min})^{S}=:\tilde{\delta}$ (100)
Let $X_{s,s^{\prime}}$ be a random variable that represents the hitting time of $s^{\prime}$ starting from $s$, i.e. it follows the conditional distribution of $\min\{t>0:s_{t}=s^{\prime}\}$ given $s_{0}=s$. Following the same derivation of Equation (1.18) of Levin and Peres (2017), we have
$\mathbb{P}_{\theta^{*}}(X_{s,s^{\prime}}>kS)\leq(1-\tilde{\delta})^{k}\quad\forall k\in\mathbb{N}$ (101)
Therefore, for any $u>0$,
$\mathbb{P}(X_{s,s^{\prime}}>u)\leq(1-\tilde{\delta})^{\lfloor u/S\rfloor}\leq(1-\tilde{\delta})^{\frac{u}{S}-1}$ (102)
Set $\lambda=-\frac{1}{2S}\log(1-\tilde{\delta})>0$, we have
$\mathbb{E}[\exp(\lambda X_{s,s^{\prime}})]$ $=\int_{0}^{\infty}\mathbb{P}(\exp(\lambda X_{s,s^{\prime}})>x)\mathrm{d}x$ (103)
$=1+\int_{1}^{\infty}\mathbb{P}(\exp(\lambda X_{s,s^{\prime}})>x)\mathrm{d}x$ (104)
$=1+\lambda\int_{0}^{\infty}\mathbb{P}(\exp(\lambda X_{s,s^{\prime}})>\exp(\lambda u))\exp(\lambda u)\mathrm{d}u$ (105)
$\leq 1+\lambda\int_{0}^{\infty}(1-\tilde{\delta})^{\frac{u}{S}-1}\exp(\lambda u)\mathrm{d}u$ (106)
$=1+\lambda\int_{0}^{\infty}(1-\tilde{\delta})^{\frac{u}{2S}-1}\mathrm{d}u=:\exp(\tilde{\lambda}).$ (107)
Next, we establish that with high probability, each state must be visited at least $c_{1}N$ times in an $N$-step trajectory for some constant $c_{1}>0$. Let $\bar{N}(s):=\sum_{t=0}^{N-1}\bm{1}_{\{\bar{s}_{t}=s\}}$ be the number of visits to state $s$ in the offline data (excluding the final state $\bar{s}_{N}$). Let $X_{\bar{s}_{1},s;0},X_{s,s;1},X_{s,s;2},X_{s,s;3},\cdots$ be mutually independent random variables representing hitting times. For any $K\in\mathbb{N}$ we have
$\mathbb{P}_{\theta^{*}}(\bar{N}(s)<K)$ $\leq\mathbb{P}_{\theta^{*}}\left(X_{\bar{s}_{1},s;0}+\sum_{k=1}^{K-1}X_{s,s;k}\geq N\right)$ (108)
$=\mathbb{P}_{\theta^{*}}\left(\exp\left(\lambda X_{\bar{s}_{1},s;0}+\lambda\sum_{k=1}^{K-1}X_{s,s;k}\right)\geq\exp(\lambda N)\right)$ (109)
$\leq\exp(-\lambda N)\mathbb{E}\left[\exp\left(\lambda X_{\bar{s}_{1},s;0}+\lambda\sum_{k=1}^{K-1}X_{s,s;k}\right)\right]$ (110)
$\leq\exp(-\lambda N)[\exp(\tilde{\lambda})]^{K}=\exp(-(\lambda N-\tilde{\lambda}K)).$ (111)
Set $K=\lfloor\frac{\lambda}{2\tilde{\lambda}}N\rfloor$, we have $\mathbb{P}_{\theta^{*}}(\bar{N}(s)<K)\leq\exp(-\lambda N/2)$.
Fix $s\in\mathcal{S}$ and let $a_{s}^{*}:=\pi^{*}(s)$. Under Assumption 3.6 we have
$\pi^{\beta}(a_{s}^{*}|s)$ $=\dfrac{\exp(\beta q(s,a_{s}^{*};\theta^{*}))}{\sum_{a\in\mathcal{A}}\exp(\beta q(s,a;\theta^{*}))}$ (112)
$=\dfrac{1}{\sum_{a\in\mathcal{A}}\exp[\beta(q(s,a;\theta^{*})-q(s,a_{s}^{*};\theta^{*}))]}$ (113)
$\geq\dfrac{1}{1+(A-1)\exp(-\beta\Delta)}.$ (114)
For the rest of the proof, we assume that $\beta\geq\log(2(A-1))/\Delta$. Then $\pi^{\beta}(a_{s}^{*}|s)\geq\frac{2}{3}$.
For $L\in\mathbb{N}$, define $\mathcal{F}_{L}(s)$ to be the event that in the first $L$ visits to the state $s$, the action $\pi^{*}(s)$ is chosen $\leq L/2$ times. We have
$\mathbb{P}_{\theta^{*}}(\mathcal{F}_{L})\leq\mathbb{P}(\mathcal{B}(L,2/3)\leq L/2)\leq\exp\left(-\frac{L}{18}\right)$ (115)
where $\mathcal{B}(L,p)$ is a binomial random variable with parameters $L$ and $p$. The last inequality in (115) can be established with Hoeffding’s inequality.
Now, consider a construction of $\hat{\pi}^{*}$ as follows: For each $s\in\mathcal{S}$ and $a\in\mathcal{A}$, let $\bar{N}(s,a):=\sum_{t=1}^{N}\bm{1}_{\{\bar{s}_{t}=s,\bar{a}_{t}=a\}}$. Define $\hat{\pi}^{*}(s)=\operatorname*{arg\,max}_{a\in\mathcal{A}}\bar{N}(s,a)$ (ties are broken following a fixed ordering on $\mathcal{A}$). Now we have
$\mathbb{P}_{\theta^{*}}(\hat{\pi}^{*}(s)\neq\pi^{*}(s))$ $\leq\mathbb{P}(\bar{N}(s,a_{s}^{*})\leq\bar{N}(s)/2)$ (116)
$\leq\mathbb{P}_{\theta^{*}}(\bar{N}(s)<K)+\mathbb{P}_{\theta^{*}}(\bar{N}(s)\geq K,\bar{N}(s,a_{s}^{*})\leq\bar{N}(s)/2)$ (117)
$\leq\mathbb{P}_{\theta^{*}}(\bar{N}(s)<K)+\sum_{L=K}^{N}\mathbb{P}_{\theta^{*}}(\mathcal{F}_{L})$ (118)
$\leq\exp\left(-\dfrac{\lambda N}{2}\right)+N\exp\left(-\frac{K}{18}\right)\leq\left(1+e^{1/18}N\right)\exp(-cN)$ (119)
where $c:=\min\{\frac{\lambda}{2},\frac{\lambda}{36\tilde{\lambda}}\}$.
Note that $c$ is a function of $(S,A,\bar{v},\Delta,\delta,\beta)$, in particular, it is independent of $\theta^{*}$. Therefore we conclude that
$\mathbb{P}(\hat{\pi}^{*}\neq\pi^{*})\leq\sum_{s\in\mathcal{S}}\mathbb{P}(\hat{\pi}^{*}(s)\neq\pi^{*}(s))\leq S(1+1.06N)\exp(-cN)$ (120)

### v2

@@
 This requires a novel prior-dependent regret analysis of Bayesian online learning algorithms for the infinite horizon setting.
-We then propose an approximate Informed RLSVI algorithm that we can interpret as performing imitation learning with the offline dataset, and then performing online learning.
+We then propose the Informed RLSVI algorithm to efficiently approximate the iPSRL algorithm.
 1 Introduction
@@
 However, determining whether the dataset meets these conditions can be a complicated problem itself, hence making these results impractical (Kumar et al., 2020; Nair et al., 2020; Argenson and Dulac-Arnold, 2020; Levine et al., 2020; Kostrikov et al., 2021; Wagenmaker and Pacchiano, 2022).
+There have also been a few experimental work on leveraging offline data for online learning (Zheng et al., 2023; Feng et al., 2023; Hu et al., 2023).
 On the other hand, there have been many works on online reinforcement learning (See Russo et al.
@@
 In (Schrittwieser et al., 2021; Uehara and Sun, 2021; Xie et al., 2021b; Agarwal and Zhang, 2022; Song et al., 2022; Fang et al., 2022; Wan et al., 2022; Ball et al., 2023), the authors combined offline RL with limited online policy fine-tuning to minimize the simple regret.
+Contributions.
+(i) We introduced the informed PSRL algorithm for infinite horizon MDPs that incorporates prior data and takes deterministic episode lengths as input (compared to Ouyang et al.
+(2017) where the episode schedule is fixed and dependent on visitation counts).
+(ii) We develop an optimal policy estimation-error based technique that differs from the existing information-ratio based techniques (Russo and Van Roy, 2016; Hao et al., 2023) for prior-dependent regret analysis.
+(iii) We provide a novel prior-dependent regret bound for learning infinite horizon MDPs that scales polynomially with all relevant parameters.
+(iv) We introduced the iRLSVI algorithm to approximate the informed PSRL algorithm.
+The algorithm can be seen as bridging online learning and imitation learning.
+Notations.
+We use script letters for sets.
+For a set $\mathcal{X}$, we use $\Delta(\mathcal{X})$ to represent the set of probability measures on $\mathcal{X}$.
+$\mathbb{P},\mathbb{E}$ stand for probability and expectation respectively.
+We use the standard big-$O$ and big-$\tilde{O}$ notations to hide absolute constants and/or logarithmic terms.
 2 Preliminaries
@@
 (2017) for more discussions.
-For a communicating MDP $\mathcal{M}(\theta)$, we use $\bar{r}^{*}(\theta),v(\cdot;\theta)$, and $q(\cdot,\cdot;\theta)$ to denote their corresponding $\bar{r}^{*},v^{*},$ and $q^{*}$ functions, as defined in (2), (3) and (5) respectively.
+For a communicating MDP $\mathcal{M}(\theta)$, we use $\bar{r}^{*}(\theta),v(\cdot;\theta)$, and $q(\cdot,\cdot;\theta)$ to denote their corresponding $\bar{r}^{*},v,$ and $q$ functions, as defined in (2), (3) and (5) respectively.
 We also make Assumption 2.2, which is the same as that in Ouyang et al.
@@
 There exists $\bar{v}\in\mathbb{R}_{+}$ such that $\max_{s\in\mathcal{S}}v(s;\theta)-\min_{s\in\mathcal{S}}v(s;\theta)\leq\bar{v}$ for all $\theta\in\varTheta$.
+Remark 2.3.
+Note that we can replace the “for all $\theta\in\varTheta$” statements in Assumption 2.1 and 2.2 with “for $\mu_{0}$-almost all $\theta\in\varTheta$” (a.k.a.
+holds for $\theta^{*}$ with probability 1) without loss of generality: If certain $\theta\in\varTheta$ does not satisfy either assumption we can remove it from the parameter set and it will affect neither our algorithm nor the analysis.
+[^1] Under certain mild measurability assumptions.
 Under the above assumptions we consider the canonical settings of tabular MDPs as well as linear value function generalization for non-tabular MDPs, i.e., the Q-value function is representable in a subspace spanned by a given set of features.
@@
 Then, we consider a general model wherein the expert only knows a noisy version of the optimal $q$-function.
-The expert uses the policy (8), except that $q^{*}$ is replaced by $\tilde{q}$, the expert’s perceived state-action value function.
-The vector $\tilde{q}$ follows Gaussian distribution with mean $q^{*}$ and covariance $\mathbb{I}/\lambda^{2}$, where $\lambda>0$ is the knowledgeability parameter.
+The expert uses the policy (8), except that $q$ is replaced by $\tilde{q}$, the expert’s perceived state-action value function.
+The vector $\tilde{q}$ follows Gaussian distribution with mean $q$ and covariance $\mathbb{I}/\lambda^{2}$, where $\lambda>0$ is the knowledgeability parameter.
 The two parameters $(\beta,\lambda)$ together will be referred to as the competence of the expert.
@@
 The learning agent knows the form of the expert’s policy and the knowledgeability parameter $\lambda$ but may not know the deliberateness parameter $\beta$.
-When $\beta$ is unknown to the learning agent, we will assume an independent exponential prior for the sake of analytical simplicity, $f_{2}(\beta(s))=\lambda_{2}\exp(-\lambda_{2}\beta(s))$ over $\beta(s)$ where $\lambda_{2}>0$ is the same for all $s$.
+For the sake of simplicity, when $\beta$ is unknown to the learning agent, we assume that $\beta(s)=\beta$ is an exponential random variables with probability density function $f_{2}(x)=\lambda_{2}\exp(-\lambda_{2}x)\bm{1}_{\{x>0\}}$ where $\lambda_{2}>0$.
 3 The Infinite-horizon Informed PSRL Algorithm for Average MDPs
@@
 can be used in MDPSolve.
+A similar MDP solving oracle is also needed for Ouyang et al.
+(2017).
 The inf-iPSRL algorithm divides the learning process into episodes, with possibly different lengths.
@@
 $\mu_{k+1}(\theta^{*}\in\cdot)\propto\int_{\theta\in\cdot}\prod_{t=t_{k}}^{t_{k+1}-1}\theta(s_{t+1}|s_{t},a_{t})\mathrm{d}\mu_{k}(\theta).$ (11)
-Algorithm 1 inf-iPSRL Algorithm
+Algorithm 1 inf-iPSRL
   Input: Prior $\mu^{0}\in\Delta(\varTheta)$; Offline data $\mathcal{D}_{0}$; Episode schedule $\{T_{k}\}_{k=1}^{K}$
@@
 $\mathbb{P}(\tilde{\pi}^{1}\neq\pi^{*})\leq\mathbb{P}(\hat{\pi}^{*}\neq\tilde{\pi}^{1})+\mathbb{P}(\hat{\pi}^{*}\neq\pi^{*})=2\mathbb{P}(\hat{\pi}^{*}\neq\pi^{*})$
+Remark 3.6.
+When $\varepsilon=\Theta(1)$, the regret bound in Corollary 3.4 has an additional factor of $\sqrt{SA}$ comparing to the $\tilde{O}(\bar{v}\sqrt{SAT})$ regret bound for (uninformed) learning in communicating MDPs given by Agrawal and Jia (2017).
+However, the advantage of our prior-dependent bound is for small $\varepsilon$.
+The next section shows that under certain assumptions, $\varepsilon$ can be exponentially small in the amount of prior data.
+We leave the improvement of the prior-dependent upper bound’s dependency on $S$ and $A$ as future work.
 3.2 Bounding the Estimation Error of Optimal Policy
@@
 4.1 The Informed RLSVI Algorithm for Average-reward MDPs
-The inf-iPSRL algorithm introduced in the previous section assumes that the posterior updates can be done exactly.
-In practice, the posterior update in (9) is challenging due to the loss of conjugacy while using the Bayes rule.
-Thus, we must find a computationally efficient way to do approximate posterior updates (and obtain samples from it) to enable practical implementation.
-Hence, we propose a novel approach based on Bayesian bootstrapping to obtain approximate posterior samples.
-The key idea is to perturb the loss function for the maximum a posterior (MAP) estimate and use the point estimate as a surrogate for the exact posterior sample.
-Note that in the ensuing, we regard $\beta$ as also unknown to the learning agent (and $\lambda=\infty$ for simplicity).
-Thus, the learning agent must form a belief over both $\theta$ and $\beta$ via a joint posterior distribution conditioned on the offline dataset $\mathcal{D}_{0}$ and the online data at time $t$, $\mathcal{H}_{t}$.
-We denote the prior pdf over $\theta$ by $f(\cdot)$ and prior pdf over $\beta$ by $f_{2}(\cdot)$.
-Consider the average-reward Bellman optimality eq.:
-$\bar{r}^{*}+q^{*}(s,a;\theta)=r(s,a)+\mathbb{E}[\max_{b}q^{*}(s^{\prime},b;\theta)].$
-For the sake of compact notation, we denote $q^{*}(s,a;\theta)$ as $q^{\theta}_{t}(s,a)$ in this section. Now, consider the offline dataset,
-$\mathcal{D}_{0}=\{(s_{k},a_{k},\check{s}_{k},r_{k})_{k=0}^{N-1}\}$
-and let $\theta$ denote the parameterization of the relative q-value function. We introduce the temporal difference error ${\mathcal{E}}_{k}$ (parameterized by a given $q^{\theta}$),
-$\mathcal{E}_{k}(q^{\theta}):=\left(r_{k}+\max_{b}q^{\theta}(\check{s}_{k},b)-q^{\theta}(s_{k},a_{k})-\bar{r}_{k}\right),$
-and define recursively,
-$\bar{r}_{k+1}:=(1-\eta_{k})\bar{r}_{k}+\eta_{k}r_{k},~{}~{}\bar{r}_{0}:=r_{0}.$
-We use this to construct a parameterized offline dataset,
-$\mathcal{D}_{0}(q^{\theta})=\{(s_{k},a_{k},\check{s}_{k},\bar{r}_{k},\mathcal{E}_{k}(q^{\theta}))_{k=0}^{N-1}\}.$
-A parameterized online dataset at time $t$, $\mathcal{H}_{t}(q^{\theta})$ can be similarly defined. To ease notation, we will regard the time $t$ during the online phase as the $(N+t)$ time. Thus,
-$\mathcal{H}_{t}(q^{\theta})=\{(s_{k},a_{k},\check{s}_{k},\bar{r}_{k},\mathcal{E}_{k}(q^{\theta}))_{k=N}^{N+t}\}\},$
-the online dataset available at time $t$.
-Note that $q^{\theta}$ is to be regarded as a parameter. Now, at time $t$, we would like to obtain a MAP estimate for $(\theta,\beta)$ by solving the following:
-$\begin{split}\textbf{MAP:}~{}~{}~{}\arg\max_{\theta,\beta}~{}&\log P(\mathcal{H}_{t}(q^{\theta})|\mathcal{D}_{0}(q^{\theta}),\theta,\beta)+\log P(\mathcal{D}_{0}(q^{\theta})|\theta,\beta)+\log f(\theta)+\log f_{2}(\beta).\end{split}$ (21)
-Denote a perturbed version of the $q^{\theta}$-parameterized offline dataset by
-$\tilde{\mathcal{D}}_{0}(q^{\theta})=\{(s_{k},\tilde{a}_{k},\check{s}_{k},\bar{r}_{k},\tilde{\mathcal{E}}_{k})_{k=0}^{N-1}\}$
-where we add the following random perturbations: (i) actions have perturbation $w_{l}^{h}\sim\exp(1)$, (ii) rewards have perturbations $z_{k}\sim\mathcal{N}(0,\sigma^{2})$, and (iii) the prior $\tilde{\theta}\sim\mathcal{N}(0,\Sigma_{0})$.
-Note that the first and second terms involving $\mathcal{H}_{t}$ and $\mathcal{D}_{0}$ in (21) are independent of $\beta$ when conditioned on the actions.
-Thus, we have a sum of log-likelihood of TD error, transition and action as follows:
-$\begin{split}&\quad~{}\log P(\tilde{\mathcal{D}}_{0}(q^{\theta})|q^{\theta})\\ &=\sum_{k=0}^{N-1}\left(\log P(\tilde{\mathcal{E}}_{k}|\check{s}_{k},a_{k},s_{k},\bar{r}_{k},q^{\theta})+\log P(\check{s}_{k}|a_{k},s_{k},q^{\theta})+\log P(a_{k}|s_{k},q^{\theta})\right)\\ &\leq\sum_{k=0}^{N-1}\left(\log P(\tilde{\mathcal{E}}_{k}|\check{s}_{k},a_{k},s_{k},\bar{r}_{k},q^{\theta})+\log\pi^{\beta}(a_{k}|s_{k},q^{\theta})\right).\end{split}$
-By ignoring the log-likelihood of the transition term (akin to optimizing an upper bound on the negative loss function), we are actually being optimistic.
-For the terms in the upper bound above, under the random perturbations assumed above, we have
-$\begin{split}&\log P(\tilde{\mathcal{E}}_{k}|\check{s}_{k},a_{k},s_{k},\bar{r}_{k},q^{\theta})=-\frac{1}{2}(r_{k}+z_{k}+\max_{b}q^{\theta}(\check{s}_{k},b)-q^{\theta}(s_{k},a_{k})-\bar{r}_{k})^{2}+\text{constant}\end{split}$
-and
-$\begin{split}\log\pi^{\beta}(a_{k}|s_{k},q^{\theta}_{t})&=w_{k}\left(\beta q^{\theta}(s_{k},a_{k})-\log\sum_{b}\exp\left(\beta q^{\theta}(s_{k},b)\right)\right).\end{split}$
-Now, denote a perturbed version of the $q^{\theta}$-parameterized online dataset by
-~$\tilde{\mathcal{H}}_{t}(q^{\theta})=\{(s_{k},a_{k},\check{s}_{k},\bar{r}_{k},\tilde{\mathcal{E}}_{k})_{k=N:N+t}\},$
-and thus similar to before, we have
-$\begin{split}&\quad~{}\log P(\tilde{\mathcal{H}}_{t}(q^{\theta})|\tilde{\mathcal{D}}_{0}(q^{\theta}),q^{\theta})\\ &=\sum_{k=N}^{N+t}\left(\log P(\tilde{\mathcal{E}}_{k}(q^{\theta})|\check{s}_{k},a_{k},s_{k},\bar{r}_{k},q^{\theta})+\log P(\check{s}_{k}|a_{k},s_{k},q^{\theta})\right),\\ &\leq\sum_{k=N}^{N+t}\log P(\tilde{\mathcal{E}}_{k}|\check{s}_{k},a_{k},s_{k},\bar{r}_{k},q^{\theta}),\\ \end{split}$
-where we again ignored the transition term to obtain an optimistic upper bound.
-Given the random perturbations above, we have
-$\begin{split}&\quad~{}\log P(\tilde{\mathcal{E}}_{k}(q^{\theta})|\check{s}_{k},a_{k},s_{k},\bar{r}_{k},q^{\theta})\\ &=-\frac{1}{2}\left(r_{k}+z_{k}+\max_{b}q^{\theta}(\check{s}_{k},b)-q^{\theta}(s_{k},a_{k})-\bar{r}_{k}\right)^{2}+\text{constant}.\end{split}$
-The prior over $\beta$, $f_{2}(\beta)$ is assumed to be an exponential pdf $\lambda_{2}\exp(-\lambda_{2}\beta),\beta\geq 0$, while that over $\theta$ is assumed Gaussian.
-Thus, putting it all together, we get the following optimistic loss function (to minimize over $\theta$ and $\beta$),
-~$\begin{split}\tilde{\mathcal{L}}(\theta,\beta)&=\frac{1}{2\sigma^{2}}\sum_{k=0}^{N+T-1}\left(r_{k}+z_{k}+\max_{b}q^{\theta}(\check{s}_{k},b)-q^{\theta}(s_{k},a_{k})-\bar{r}_{k}\right)^{2}\\ &-\sum_{t=0}^{T-1}w_{k}\left(\beta q^{\theta}(s_{t},a_{t})-\log\sum_{b}\exp\left(\beta q^{\theta}(s_{t},b)\right)\right)+\frac{1}{2}(\theta-\tilde{\theta})^{\top}\Sigma_{0}(\theta-\tilde{\theta})+\lambda_{2}\beta.\end{split}$ (22)
-The above loss function is difficult to optimize in general due to the $\max$ operation, and the $Q$-value function in general having a nonlinear form.
+In the previous section, we considered the inf-iPSRL algorithm where the learning agent performs an posterior update at each episode.
+However, computing exact posterior updates can be challenging due to the loss of conjugacy when applying Bayes rule in (9).
+In this section, we propose a novel Baysian bootstrapping inspired approach to perform approximate posterior updates, where we use the maximum a posterior (MAP) estimator based on a perturbed loss function as a surrogate for samples from exact posterior distributions.
+In this section, we consider a different setting as the previous section: the expert deliberateness parameter $\beta$ is assumed to be unknown to the learning agent.
+Instead, $\beta(s)=\beta$ is an exponential random variable with probability density function $f_{2}$.
+At each time in the algorithm, the learning agent form a joint belief over $(\theta,\beta)$ based on all the offline and online data.
+The prior distribution of $(\theta,\beta)$ is given by the density function $f\times f_{2}$.
+To unify notations for offline and online data, define $\check{s}_{k}$ to be $\bar{s}_{k}$ if $k<N$ and $s_{k-N}$ otherwise).
+Define $\check{a}_{k}$ similarly.
+To simplify notations, write $q(s,a;\theta)$, defined in (5), as $q^{\theta}_{t}(s,a)$.
+Write $r_{k}=r(\check{s}_{k},\check{a}_{k})$.
+Consider the average-reward Bellman optimality equation
+$\bar{r}^{*}+q(s,a;\theta)=r(s,a)+\mathbb{E}_{s^{\prime}\sim\theta(s,a)}\left[\max_{a^{\prime}\in\mathcal{A}}q(s^{\prime},a^{\prime};\theta)\right].$ (21)
+Let $\tilde{r}_{k}$ be defined recursively via
+$\tilde{r}_{k+1}:=(1-\eta_{k})\tilde{r}_{k}+\eta_{k}r_{k},~{}~{}\tilde{r}_{0}:=r_{0}.$
+Define the temporal difference error ${\mathcal{E}}_{k}$ (parameterized by a given $q^{\theta}$),
+$\mathcal{E}_{k}(q^{\theta}):=\left(r_{k}+\max_{a^{\prime}\in\mathcal{A}}q^{\theta}(\check{s}_{k}^{\prime},a^{\prime})-q^{\theta}(\check{s}_{k},\check{a}_{k})-\tilde{r}_{k}\right),$
+where $\check{s}_{k}^{\prime}=\check{s}_{k+1}$ for all $k$ except when $k=N-1$, in which case $\check{s}_{k}^{\prime}=\bar{s}_{N}$.
+Here $\tilde{r}_{k}$ is used as a bootstrap approximation for the true optimal average reward $\bar{r}^{*}$.
+Therefore, conditioning on $(\theta,\check{s}_{k},\check{a}_{k})$, $\mathcal{E}_{k}(q^{\theta})$ is approximately zero-mean.
+Now, for the convenience of describing the algorithm, define the parameterized dataset, which contains offline and online data up to (offline/online combined) time $t$:
+$\mathcal{H}_{t}(q^{\theta}):=\{(\mathbf{p}_{k})_{k=0}^{t}\}:=\{(\check{s}_{k},\check{a}_{k},\mathcal{E}_{k}(q^{\theta}))_{k=0}^{t}\}\},$ (22)
+At time $t$, consider the following simplistic way of viewing the parameterized dataset: (i) Each $\mathbf{p}_{k}$ is a separate data point where $\check{s}_{k}$ is viewed as a given context.
+(ii) Conditioning on $(\theta,\check{s}_{k},\check{a}_{k})$, the temporal difference error $\mathcal{E}_{k}(q^{\theta})\sim\mathcal{N}(0,\sigma^{2})$.
+(iii) The actions in the offline part of the dataset contains additional information (compared to the online part) since the actions were generated through policy $\pi^{\beta}(\cdot|\cdot,q^{\theta})$.
+Otherwise the actions are also given context.
+In this view, the maximum a posterior (MAP) estimate of the parameters $(\theta,\beta)$ at (online) time $t$ is given by $\arg\min_{\theta,\beta}\mathcal{L}(\theta,\beta)$ where
+$\mathcal{L}(\theta,\beta)$ $=-\sum_{k=0}^{N+t-1}\log\tilde{\mathbb{P}}_{\theta}(\mathcal{E}_{k}(q^{\theta})|\check{s}_{k},\check{a}_{k})-\sum_{k=0}^{N-1}\log\pi^{\beta}(\check{a}_{k}|\check{s}_{k};q^{\theta})-\log f(\theta)-\log f_{2}(\beta)$ (23)
+$=\underbrace{\frac{1}{2\sigma^{2}}\sum_{k=0}^{N+t-1}\left(r_{k}+\max_{a^{\prime}\in\mathcal{A}}q^{\theta}(\check{s}_{k},a^{\prime})-q^{\theta}(s_{k},a_{k})-\tilde{r}_{k}\right)^{2}}_{\text{Value Iteration Loss}}$ (24)
+$\underbrace{-\sum_{k=0}^{N-1}w_{k}\left(\beta q^{\theta}(\bar{s}_{k},\bar{a}_{k})-\log\sum_{a^{\prime}\in\mathcal{A}}\exp\left(\beta q^{\theta}(\bar{s}_{k},a^{\prime})\right)\right)}_{\text{Imitation Loss}}+\underbrace{\frac{1}{2}\theta^{\top}\Sigma_{0}\theta+\lambda_{2}\beta}_{\text{Prior Loss}}+~{}\text{constant}.$ (25)
+Now we add the following random perturbations to the loss function: (i) each temporal difference error term is perturbed by $z_{k}\sim\mathcal{N}(0,\sigma^{2})$, (ii) each imitation loss term is weighted by a random $w_{k}\sim\exp(1)$, and (iii) $\theta$ in the prior loss term is perturbed by $\tilde{\theta}\sim\mathcal{N}(0,\Sigma_{0})$.
+Combining the above, we get the following randomized loss function,
+~$\begin{split}\tilde{\mathcal{L}}(\theta,\beta)&=\underbrace{\frac{1}{2\sigma^{2}}\sum_{k=0}^{N+t-1}\left(r_{k}+z_{k}+\max_{a^{\prime}\in\mathcal{A}}q^{\theta}(\check{s}_{k},a^{\prime})-q^{\theta}(s_{k},a_{k})-\tilde{r}_{k}\right)^{2}}_{\text{Randomized Value Iteration Loss}}\\ &\underbrace{-\sum_{t=0}^{N-1}w_{k}\left(\beta q^{\theta}(s_{t},a_{t})-\log\sum_{a^{\prime}\in\mathcal{A}}\exp\left(\beta q^{\theta}(s_{t},a^{\prime})\right)\right)}_{\text{Randomized Imitation Loss}}+\underbrace{\frac{1}{2}(\theta-\tilde{\theta})^{\top}\Sigma_{0}(\theta-\tilde{\theta})+\lambda_{2}\beta}_{\text{Randomized Prior Loss}}.\end{split}$ (26)
+Algorithm 2 inf-iRLSVI
+  Input: Prior $\mu^{0}\in\Delta(\varTheta)$; Offline data $\mathcal{D}_{0}$; Episode schedule $\{T_{k}\}_{k=1}^{K}$
+  $t_{1}\leftarrow 0$
+  for $k=1$ to $K$ do
+     Construct ~$\tilde{\mathcal{L}}(\theta,\beta)$ as described above
+     Solve $(\tilde{\theta}^{k},\tilde{\beta}^{k})=\arg\min_{\theta,\beta}\tilde{\mathcal{L}}(\theta,\beta)$
+     $\tilde{\pi}^{k}\leftarrow\texttt{MDPSolve}(\tilde{\theta}^{k})$
+     $t_{k+1}\leftarrow t_{k}+T_{k}$
+     for $t=t_{k}$ to $t_{k+1}-1$ do
+        Take action $a_{t}=\tilde{\pi}^{k}(s_{t})$
+     end for
+  end for
 Remark 4.1.
-Note that the loss function in (22) can be hard to jointly optimize over $\theta$ and $\beta$.
-In particular, estimates of $\beta$ can be quite noisy when $\beta$ is large, and the near-optimal expert policy only covers the state-action space partially.
-Thus, we consider other methods of estimating $\beta$ that are more robust, which can then be plugged into the loss function in (22).
-Specifically, we could simply look at the entropy of the empirical distribution of the actions in the offline dataset.
-Suppose the empirical distribution of $\{\bar{a}_{0}^{l},\ldots\bar{a}_{k}\}_{l=1}^{L}$ is $\mu_{A}$.
-Then we use $c_{0}/\mathcal{H}(\mu_{A})$ as an estimation for $\beta$, where $c_{0}>0$ is a hyperparameter.
-The intuition is that for smaller $\beta$, the offline actions tend to be more uniform and thus the entropy will be large.
-This is an unsupervised approach and agnostic to specific offline data generation process.
+The MAP estimation of $\beta$ can be quite noisy when $\beta$ is large due to the fact that the expert takes non-optimal actions with very small probability.
+To resolve this, we may consider an entropy-based estimate of $\beta$: The estimate is set to equal $c_{0}/\hat{H}$, where $\hat{H}$ is the conditional entropy of $\hat{a}$ given $\hat{s}$ if $(\hat{s},\hat{a})$ follows the empirical distribution given by visitation counts in the offline data.
 Remark 4.2.
-In the loss function in (22), the parameter $\theta$ appears inside the $\max$ operation.
-Thus, it can be quite difficult to optimize over $\beta$.
-Since the loss function is typically optimized via an iterative algorithm such as a gradient descent method, a simple and scalable solution that works well in practice is to use the parameter estimate $\theta$ from the previous iteration inside the $\max$ operation, and thus optimize over $\theta$ only in the other terms.
-4.2 inf-iRLSVI  bridges Online RL and Imitation Learning
-In the previous subsection, we derived inf-iRLSVI, a Bayesian-bootstrapped algorithm.
-We now present interpretation of the algorithm as bridging online RL (via commonality with the RLSVI algorithm (Osband et al., 2016) and imitation learning, and hence a way for its generalization.
-Consider the RLSVI algorithm for online reinforcement learning as introduced in Osband et al.
-(2019).
-It draws its inspiration from the posterior sampling principle for online learning, and has excellent cumulative regret performance.
-The RLSVI algorithm, that uses all of the data available at the end of episode $t$, including any offline dataset involves minimizing the corresponding loss function at each time step:
-~$\begin{split}\tilde{\mathcal{L}}_{\text{RLSVI}}(\theta)&=\frac{1}{2\sigma^{2}}\sum_{k=1}^{L+t}\sum_{h=0}^{H-1}\left(r_{k}+\max_{b}Q^{\theta}_{t+1}(\check{s}_{k},b)-q^{\theta}_{t}(s_{k},a_{k})\right)^{2}\\ &+\frac{1}{2}({\theta_{0:H}}-{\tilde{\theta}_{0:H}})^{\top}\Sigma_{0}({\theta_{0:H}}-{\tilde{\theta}_{0:H}}).\end{split}$
-Now, let us consider an imitation learning setting.
-Let $\tau_{l}=(s_{k},a_{k},\check{s}_{k})_{h=0}^{H-1}$ be the trajectory of the $l$th episode.
-Let $\hat{\pi}_{t}(a|s)$ denote the empirical estimate of probability of taking action $a$ in state $s$ at time $h$, i.e., an empirical estimate of the expert’s randomized policy.
-Let $p(\tau)$ denote the probability of observing the trajectory under the policy $\hat{\pi}$.
-Let $\pi_{t}^{\beta,\theta}(\cdot|s)$ denote the parametric representation of the policy used by the expert.
-And let $p^{\beta,\theta}(\tau)$ denote the probability of observing the trajectory $\tau$ under the policy ${\pi}^{\beta,\theta}$.
-Then, the loss function corresponding to the KL divergence between $\Pi_{l=1}^{L}p(\tau_{l})$ and $\Pi_{l=1}^{L}p^{\beta,\theta}(\tau_{l})$ is given by
-~$\tilde{\mathcal{L}}_{\text{IL}}(\beta,\theta)$ $=D_{KL}\left(\Pi_{l=1}^{L}p(\tau_{l})||\Pi_{l=1}^{L}p^{\beta,\theta}(\tau_{l})\right)=\int\Pi_{l=1}^{L}p(\tau_{l})\log\frac{\Pi_{l=1}^{L}p(\tau_{l})}{\Pi_{l=1}^{L}p^{\beta}(\tau_{l})}$ (23)
-$=\sum_{l=1}^{L}\int p(\tau_{l})\log\frac{p(\tau_{l})}{p^{\beta,\theta}(\tau_{l})},=\sum_{l=1}^{L}\sum_{h=0}^{H-1}\log\frac{\hat{\pi}_{t}(a_{k}|s_{k})}{\pi^{\beta,\theta}_{t}(a_{k}|s_{k})}$ (24)
-$=\sum_{l=1}^{L}\sum_{h=0}^{H-1}[\log\hat{\pi}_{t}(a_{k}|s_{k})-\log\pi_{t}^{\beta,\theta}(a_{k}|s_{k})]$ (25)
-$=-\sum_{l=1}^{L}\sum_{h=0}^{H-1}\left(\beta q^{\theta}_{t}(s_{k},a_{k})-\log\sum_{b}\exp\left(\beta q^{\theta}_{t}(s_{k},b)\right)\right)+\text{constant}.$ (26)
-Remark 4.3.
-(i) The loss function ~$\tilde{\mathcal{L}}_{\text{IL}}(\beta,\theta)$ is the same as the second (action-likelihood) term in (22) while the loss function ~$\tilde{\mathcal{L}}_{\text{RLSVI}}(\theta)$ is the same as the first and third terms there (except for perturbation) and minus the $\lambda_{2}\beta$ term that corresponds to the prior over $\beta$.
-(ii) Note that while we used the more common KL divergence for the imitation learning loss function, use of log loss would yield the same outcome.
-Thus, the inf-iRLSVI loss function can be viewed as
-~$\tilde{\mathcal{L}}(\beta,\theta)=\tilde{\mathcal{L}}_{\text{RLSVI}}(\theta)+\tilde{\mathcal{L}}_{\text{IL}}(\beta,\theta)+\lambda_{2}\beta,$ (27)
-thus establishing that the proposed algorithm may be viewed as bridging Online RL with Imitation Learning.
-Note that the last term corresponds to the prior over $\beta$.
-If $\beta$ is known (or uniform), it will not show up in the loss function above.
-The above also suggests a possible way to generalize and obtain other online learning algorithms that can bootstrap by use of offline datasets.
-Namely, at each step, they can optimize a general loss function of the following kind:
-~$\tilde{\mathcal{L}}_{\alpha}(\beta,\theta)=\alpha\tilde{\mathcal{L}}_{\text{ORL}}(\theta)+(1-\alpha)\tilde{\mathcal{L}}_{\text{IL}}(\beta,\theta)+\lambda_{2}\beta,$ (28)
-where ~$\tilde{\mathcal{L}}_{\text{ORL}}$ is a loss function for an Online RL algorithm, ~$\tilde{\mathcal{L}}_{\text{IL}}$ is a loss function for some Imitation Learning algorithm, and factor $\alpha\in[0,1]$ provides a way to tune between emphasizing the offline imitation learning and the online reinforcement learning.
+Due to the maximum operation inside and the non-linearity of $q$-function with respect to $\theta$, optimizing (26) can be a difficult problem in general.
+For iterative optimization algorithms such as gradient descent methods, one simple and efficient solution is to replace the maximum operation with the parameter estimate of $\theta$ in the previous iteration and optimize $\theta$ over other terms.
+4.2 Informd RLSVI Bridges Online RL and Imitation Learning
+In the previous section, the loss function can be seen as sum of a few different terms.
+In this section, we present a detailed interpretation of these terms to conclude that the algorithm can be seen as bridging the online RL algorithms (Osband et al., 2019) with imitation learning.
+In (Osband et al., 2019), the author considered episodic learning of finite horizon MDPs with horizon $H$.
+The algorithm was inspired by the posterior sampling approach and has great empirical performance.
+At the end of episode $t$, the RLSVI algorithm considers randomized loss function
+~$\begin{split}\tilde{\mathcal{L}}_{\text{RLSVI}}(\theta)&=\frac{1}{2\sigma^{2}}\sum_{k=1}^{N+t}\sum_{h=0}^{H-1}\left(r_{k,h}+z_{k,h}+\max_{a^{\prime}\in\mathcal{A}}Q^{\theta}_{h+1}(\check{s}_{k,h+1},a^{\prime})-Q^{\theta}_{h}(\check{s}_{k,h},a_{k})\right)^{2}\\ &+\frac{1}{2}({\theta_{0:H}}-{\tilde{\theta}_{0:H}})^{\top}\Sigma_{0}({\theta_{0:H}}-{\tilde{\theta}_{0:H}}).\end{split}$ (27)
+where $z_{k,h}\sim\mathcal{N}(0,\sigma^{2})$ are random perturbations.
+The sum of randomized value iteration loss and randomized prior loss in (26) can be seen as an adaptation of (27) for infinite horizon MDP learning problems.
+Let $\hat{\pi}_{t}(a|s)$ denote the empirical distribution of actions at state $s$ with all the online and offline data available at time $t$.
+The KL-divergence imitation learning loss at state $s\in\mathcal{S}$ can be specified by
+$\mathcal{L}_{\text{IL}}(\beta,\theta,s)$ $=D_{KL}\left(\hat{\pi}(s)\|\pi^{\beta}(s;q^{\theta})\right)=\sum_{a}\hat{\pi}(a|s)\log\left(\dfrac{\hat{\pi}(a|s)}{\pi^{\beta}(a|s;q^{\theta})}\right)$ (28)
+$=\sum_{a}\hat{\pi}(a|s)\log\left(\dfrac{1}{\pi^{\beta}(a|s;q^{\theta})}\right)+\text{constant}$ (29)
+$=-\dfrac{1}{\bar{N}(s)}\sum_{t=0}^{N-1}\sum_{a\in\mathcal{A}}\bm{1}_{\{\bar{s}_{k}=s,\bar{a}_{k}=a\}}\log\left(\pi^{\beta}(a|s;q^{\theta})\right)+\text{constant}$ (30)
+$=-\dfrac{1}{\bar{N}(s)}\sum_{t=0}^{N-1}\bm{1}_{\{\bar{s}_{k}=s\}}\left(\beta q^{\theta}_{t}(\bar{s}_{k},\bar{a}_{k})-\log\sum_{a^{\prime}\in\mathcal{A}}\exp\left(\beta q^{\theta}_{t}(\bar{s}_{k},a^{\prime})\right)\right)+\text{constant}.$ (31)
+where $\bar{N}(s)$ is the visitation count of state $s$ in the offline data.
+Therefore, the (pre-randomization) imitation loss in (25) can be written as a weighted sum of the individual imitation loss defined above
+[…]
+Therefore, we have shown that the randomized loss function (26) can be viewed as the sum of the RLSVI loss related to online learning, and the randomized imitation loss related to imitation learning.
+Thus the informed RLSVI algorithm can be viewed as bridging online RL with imitation learning.
 5 Conclusions
@@
 The analysis involves several novel elements and depends on prior-dependent regret analysis, which is known to be quite challenging and itself a novelty.
-We then presented an approximate Bayesian online learning algorithm, inf-iRLSVI that can be seen as naturally bridging online reinforcement learning and imitation learning.
+We then presented an approximate Bayesian online learning algorithm, inf-iRLSVI, that bridges online learning with imitation learning.
 In prior work (Osband et al., 2019), such algorithms are well-known to be computationally practical in problems that need deep exploration.